### v3

# GranQ: Granular Zero-Shot Quantization with Unified Layer-Channel Awareness

Abstract
Zero-shot quantization (ZSQ) enables neural network compression without training data, which is crucial in restricted data access environments. However, existing ZSQ methods suffer from significant activation loss in low-bit environments owing to their coarse-grained scaling strategy. To address this issue, we propose GranQ, a novel ZSQ approach that leverages layer-channel awareness to minimize the quantization error. Unlike conventional layer- or channel-wise quantization, GranQ dynamically adjusts quantization granularity by considering both layer- and channel-level activation distributions. This enables fine-grained quantization while minimizing activation distortion. Additionally, we introduce vectorized activation quantization, which enables efficient parallel computation and reduces computational overhead while preserving accuracy. GranQ achieves superior performance compared with those of state-of-the-art ZSQ methods that employ quantization-aware training. With these findings, we anticipate that GranQ will inspire novel research directions beyond conventional ZSQ approaches focused on data generation and model training. The official code will be released upon completion of the review.
1 Introduction
Figure 1: Comparison of quantization strategies. Existing methods apply either per-tensor scaling with layer-wise quantization or per-channel scaling. GranQ combines per-channel scaling with per-layer iteration for more granular quantization.
Neural network compression techniques have been extensively studied to make large-scale deep learning (DL) models practically deployable. In particular, reducing model size while minimizing performance degradation is crucial for utilizing DL on edge devices (e.g., mobile phones, embedded systems, and drones). Major approaches to model compression [7, 12] include quantization [24, 30, 17], pruning [28, 16, 33, 6], knowledge distillation [21, 20, 40], and neural architecture search [45, 13]. Among these, quantization has emerged as the most actively studied technique. It serves as an effective compression method by reducing unnecessary representational ranges in the model while maintaining its performance. However, a major limitation of quantization is the need for fine-tuning or calibration to maintain the performance of the full-precision (FP) model [17]. To address this limitation, zero-shot quantization (ZSQ) (i.e., data-free quantization) has been introduced [34, 41], which allows compressed models to maintain their performance without requiring access to the original training data. Since the introduction of ZeroQ [4], studies on ZSQ have advanced in two main directions. The first direction focuses on data generation, where synthetic data are created from the FP model. The second direction focuses on effectively applying the activation distributions of the synthetic data to the quantized (Q) model. This second direction is further divided into post-training quantization (PTQ), which calibrates the activation distributions, and quantization-aware training (QAT), which fine-tunes the Q model directly. We categorize existing ZSQ methods based on these directions, as summarized in Table 1.
Data Generation
(PTQ, QAT)
Calibration
Fine-tuning
ZeroQ (CVPR 20) [4]
GDFQ (ECCV 20) [41]
DSG (CVPR 21) [43]
MixMix (CVPR 21) [31]
Qimera (NeurIPS 21) [9]
IntraQ (CVPR 22) [44]
GENIE (CVPR 23) [25]
AdaDFQ (CVPR 23) [36]
Casual-DFQ (ICCV 23) [39]
TexQ (NeurIPS 23) [5]
RIS (AAAI 24) [1]
GenQ (ECCV 24) [32]
SQuant (ICLR 22) [18]
UDFC (ICCV 23) [2]
AIT (CVPR 22) [10]
PLF (CVPR 24) [14]
SynQ (ICLR 25) [26]
GranQ (Ours)
Table 1: Categorizing zero-shot quantization algorithms. PTQ denotes post-training quantization, and QAT denotes quantization-aware training. Data generation methods typically support both PTQ and QAT, whereas violet-marked methods utilize fine-tuning for data generation and support only QAT.
First, data generation studies focus on generating high-quality data to effectively train the Q model. A common approach uses batch normalization statistics from the FP model [41]. Recent studies focused on generating data adaptable to various bit-widths [36] and using diverse generative models such as stable diffusion guided by text prompts [32]. Meanwhile, calibration (PTQ) studies aim to minimize the quantization error by calibrating the Q model without additional training [18, 2]. Finally, fine-tuning (QAT) studies focus on transferring key information from the FP model to the Q model through knowledge distillation [10, 14, 23, 26].
However, despite extensive studies, severe performance degradation in low-bit quantization remains unresolved. To address this issue, we performed an in-depth analysis of the ZSQ process, focusing on why low-bitwidth settings still suffer from performance loss even after QAT fine-tuning. Our findings reveal that quantization errors mainly stem from the loss of activation values instead of data quality or training methods. Notably, we discovered that layer-wise (per-tensor) quantization fails to accurately preserve activations in ZSQ, leading to coarse and inaccurate representations.
Based on this analysis, we introduce GranQ, a novel ZSQ approach. GranQ independently handles large activations in specific layers and channels by narrowing the range between the maximum and minimum values within each channel (Figure 1). This dynamic adjustment minimizes activation loss and preserves the original activation values by reducing quantization errors, as shown in Figure 2. The proposed method effectively handles activation loss in low-bit quantization and achieves state-of-the-art (SOTA) performance in QAT settings on the CIFAR and ImageNet datasets. Furthermore, we implement vectorized operations for granular quantization with minimal computational overhead, enabling efficient quantization on GPUs. Our contributions can be summarized as follows:
  • We perform a comprehensive analysis of quantization errors in low-bit ZSQ settings. Our findings reveal that conventional activation quantization methods (e.g., layer- and channel-wise) lead to significant activation loss. This issue stems from coarse-grained activation quantization, which fails to preserve fine-grained details.
  • We propose GranQ, a novel method that supports granular quantization and maintains computational efficiency. We introduce a fine-grained quantization mechanism, which considers every channel in each layer, to reduce activation loss. Additionally, we mitigate computational overhead using vectorized activation quantization. To the best of our knowledge, our approach is the first to address the ZSQ problem.
  • We achieve SOTA performance over existing ZSQ methods through extensive evaluation. Specifically, on the CIFAR-100 dataset, in the 3-bit quantization setting, GranQ achieves an accuracy of 62.73%, improving by 5.45% over the latest method on ResNet-20. Furthermore, on the CIFAR-10 dataset in the 5-bit quantization setting, GranQ achieves an accuracy of 94.06%, slightly exceeding the FP model performance by 0.17%.
(a) Layer Quantization
(b) GranQ (Ours)
Figure 2: Comparison between (a) layer-wise quantization and (b) the proposed GranQ method on the CIFAR-10 dataset from AdaDFQ[36]. We visualize the activation of the first layer in ResNet-20 for the first image of the first batch. $w$ and $a$ denote weight and activation, respectively. Each sub-figure presents the activation graph for 32-bit FP (left) and 3-bit quantization (right). Although layer-wise quantization significantly distorts the original activation after quantization, GranQ effectively preserves it with minimal distortion.
2 Related Work
2.1 Quantization
Quantization can reduce the representational range of pre-trained deep neural networks (DNNs) and minimize memory usage. Quantization methods are categorized into PTQ and QAT based on the timing of the range reduction.
PTQ quantizes fully trained DNNs without any further training [17, 3], offering low computational cost and simple implementation. However, because PTQ directly applies quantization to the FP model, it struggles to properly adjust the range during quantization. To overcome this, calibration is used to select optimal quantization ranges by determining scaling factors. In contrast, QAT fine-tunes the Q model with quantization operations integrated into the training process, enabling direct optimization of activations [17]. In particular, it employs the straight-through estimator [42] to estimate gradients.
However, both PTQ and QAT have the limitation of requiring access to data. PTQ needs data for calibration, whereas QAT requires it for fine-tuning. In practice, obtaining training data is often challenging. To address this problem, ZSQ has been proposed, which performs quantization without using original data [41].
2.2 Zero-shot Quantization
As summarized in Table 1, ZSQ was initially introduced by ZeroQ [4], leading to the development of various ZSQ algorithms. These studies have explored diverse approaches to improve quantization performance under data-free settings. Recent studies have focused on generating high-quality synthetic data [25, 32] and developing more effective methods to utilize them [10, 23, 26]. However, these approaches are still limited in preserving activations in low-bit settings. Therefore, a new quantization method is required to effectively minimize activation loss. In this study, we propose a novel approach to address this challenge.
3 Preliminaries and Problem Definition
In this section, we provide a new definition of the existing ZSQ problem and introduce the preliminaries. The detailed mathematical notations are listed in Table 6 in the supplementary material.
3.1 Activation Quantization
Activation quantization reduces the precision of intermediate activation values in neural networks by converting them into low-bitwidth representations [8]. This technique is often used alongside weight quantization to further compress the model. Among various activation quantization methods, linear quantization is the most widely used. It uses a normalized scaling factor to adjust the activation values, and subsequently converts them into integers within a fixed bit range [24, 15]. The overall process typically involves three stages: forward quantization, scaling parameter computation, and dequantization [17].
3.1.1 Forward Quantization
The forward quantization process involves mapping continuous floating-point values to an integer space. This process utilizes the quantization operator $Q$, and the equation is defined as follows:
$x_{q}=\mathcal{Q}(x,s,z)=\left\lfloor\frac{x}{s}+z\right\rceil$ (1)
[…]
3.1.2 Scaling Parameter Computation
In activation quantization, setting the scaling factor $s$ is crucial for mapping various activation distributions to a fixed integer range. In this study, we employ the asymmetric quantization method, which is the most commonly used for activation quantization [29]. We calculate the scaling factor $s$ and the zero-point $z$ as follows:
$s=\frac{x_{\max}-x_{\min}}{2^{b}-1},\quad z=\left\lfloor\frac{-x_{\min}}{s}\right\rceil$ (2)
Here, $x_{\max}$, $x_{\min}$ and $b$ represent the maximum and minimum values of the activation, and the number of quantization bits, respectively. The zero-point $z$ represents the offset used to adjust activation values. $x_{\max}$ and $x_{\min}$ determine the range of the activation distribution, which directly affects the computation of scaling factor $s$. Specifically, when $x_{\max}-x_{\min}$ is large, the scaling factor $s$ increases. This leads to a larger quantization step size, making it harder to preserve fine-grained activations. Conversely, when $x_{\max}-x_{\min}$ is small, the scaling factor $s$ also decreases. This allows for finer value distinctions and enhances the quality of representation.
3.1.3 Dequantization
The dequantization process converts quantized integer values back into their real-valued representation, as shown in the following equation:
$x_{\text{deq}}=s(x_{q}-z)$ (3)
In this process, the zero-point $z$ is subtracted to reverse the forward quantization, and the result is multiplied by the scaling factor $s$ to obtain the final value $x_{deq}$. The dequantization process restores quantized activation values to their original real-valued range. It is essential during model inference for converting final predictions into real values. This process is essential for converting the final prediction into real values during model inference.
Figure 3: Visualization of activation magnitudes for the first image in the first batch using ResNet-20 on CIFAR-10. The left column illustrates activations from the FP model, whereas the right column presents activations after layer-wise quantization. Significant distortions are observed in certain layers, particularly in the final layer.
3.2 Problem Definition
Existing activation quantization methods are designed based on layer-wise quantization to minimize computational cost. However, fixed scaling factors struggle to handle varying activations. This issue becomes more severe under low-bitwidth settings in ZSQ environments. The following outlines the primary challenges faced in ZSQ under low-bitwidth settings.
(Problem 1) Coarse-grained activation representation
In conventional activation quantization approaches, the scaling factor $s$ is fixed per layer or channel, mapping all activations to the same range. However, in low-bitwidth settings, this approach struggles to capture fine-grained activation variations from generated data.
(Problem 2) Severe quantization error in certain layers
During quantization, certain layers induce severe quantization errors, leading to significant activation distortions. These distortions substantially differ from the original distribution and can directly impact the performance of the model.
4 Observation and Methodology
We observe the causes of activation distortion during quantization and analyze how to mitigate them. Based on this, we introduce the proposed method known as GranQ.
4.1 Observation
ResNet-20 Layer-Q Channel-Q GranQ
$Avg.\frac{X\cdot Q}{\|X\|_{2}\|Q\|_{2}}(\uparrow)$
$Avg.\frac{\|X-Q\|_{2}}{\|X\|_{2}}(\downarrow)$
0.5111
0.3129
0.5602
0.2528
0.6835
0.1063
Table 2: Comparative analysis of averaged cosine similarity and relative error over all layers and batches. Higher cosine similarity ($\uparrow$) indicates better retention of activation information, whereas lower relative error ($\downarrow$) signifies reduced quantization distortion. Layer-Q and Channel-Q denote layer- and channel-wise quantization, respectively, whereas GranQ integrates both for improved quantization.
Layer-wise quantization effectively captures the activation distribution of each layer. However, it becomes inefficient when activation distributions within a layer exhibit significant variation along the channel axis. Particularly in ZSQ, as shown in the FP graph in Figure 2(a), the activation distributions can differ substantially across channels even within the same layer. This implies that limited representation bits are forced to capture a wide range of activation values, making fine-grained quantization difficult. It consistently appears across all layers. We analyzed this phenomenon occurring across all layers and provided a summary in Table 2. In Table 2, we present the cosine similarity $\frac{X\cdot Q}{|X|_{2}|Q|_{2}}$, along with the relative error $\frac{|X-Q|_{2}}{|X|_{2}}$. From Table 2, we identify that both layer- and channel-wise quantization fail to preserve the original activation effectively. This indicates that layer- and channel-wise quantization, which independently process activations at the layer or channel level, are not suitable for ZSQ. In contrast, GranQ, which considers both layers and channels simultaneously, achieves 1.34$\times$ higher similarity and 2.94$\times$ lower relative error compared with those of layer-wise quantization, demonstrating its effectiveness in preserving activations.
Furthermore, Figure 3 provides a visual representation of the effect of quantization on activation values across layers. In layer-quantizaiton, although the overall activations generally reflect the FP activations to some extent, certain layers exhibit fine activation values being mapped to discrete values, leading to coarse representations. These coarse representations have been identified as a major factor leading to a decline in the overall expressiveness and inference performance of the model. Further analysis of Figure 3 is provided in the supplementary material, in Figures 6 and 7.
4.2 Granular Quantization
Based on our analysis, we propose GranQ (granular layer-channel quantization), a fine-grained ZSQ approach designed to minimize computational overhead, as shown in Figure 4. GranQ combines layer and channel information using layer-channel scaling and improves efficiency through vectorized quantization.
4.2.1 Layer-Channel Scaling
We propose layer-channel scaling, which enables GranQ to assign individual scaling factors to each channel, enhancing quantization precision. This approach reduces the impact of outlier activations in specific channels on the quantization process. The layer-channel scaling is defined as follows:
$\vec{x}_{\min}$ $=\min_{h,w}(x_{l}),\quad\vec{x}_{\max}=\max_{h,w}(x_{l})$ (4)
where $\text{where }x_{l}\in\mathbb{R}^{C\times H\times W}$
[…]
Dataset
Model
(FP 32)
Bits
GDFQ
(ECCV 20)
ARC+AIT
(CVPR 22)
AdaDFQ
(CVPR 23)
TexQ
(NeurIPS 24)
AIT+RIS
(AAAI 24)
PLF
(CVPR 24)
GenQ
(ECCV 24)
AKT
(SAC 25)
SynQ
(ICLR 25)
GranQ
[…]
Table 3: Accuracy Evaluation of QAT Methods for ZSQ. $w$ and $a$ represent weight and activation, respectively. Bold values indicate the best accuracy, and underlined values denote the second-best accuracy. $\dagger$ indicates our re-implementation with the official code.
[…]
4.2.2 Vectorized Quantization
In the layer-channel scaling stage, the activation tensor of each layer is vectorized along the channel dimension. In the subsequent vectorized quantization stage, both the scaling factor $\vec{s}$ and zero-point $\vec{z}$ are also vectorized and computed as shown in Equation 5. The activation values go through quantization and dequantization (Equation 6), with all operations running in parallel across channels. During computation, scaling factor $s$ and zero-point $z$ of each channel are broadcasted to the activation values and applied element-wise.
$\vec{s}=\frac{\vec{x}_{\max}-\vec{x}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{x}_{\min}}{\vec{s}}\right\rceil$ (5)
$x_{q,l}=\left\lfloor\frac{x}{\vec{s}}+\vec{z}\right\rceil,x_{\text{deq},l}=\vec{s}(x_{q,l}-\vec{z})$ (6)
This process can be applied to every layer in the network, enabling the vectorized quantization of all activation tensors. A detailed computational process is provided in Algorithm 1, available in the supplementary material.
5 Experiments
In this section, we thoroughly evaluate the effectiveness of GranQ. Experiments are conducted on diverse benchmark datasets, with the performance compared with those of existing ZSQ methods.
5.1 Experimental Setup and Details
The experiments were conducted using widely adopted ZSQ evaluation datasets, including CIFAR-10, CIFAR-100 [27], and ImageNet (ILSVRC 2012) [11] validation datasets. For the CIFAR datasets, ResNet-20 [19] was used as the quantization model, whereas ResNet-18 [19], ResNet-50 [19], and MobileNetV2 [38] were employed for ImageNet. All experiments were conducted using the SGD optimizer [37] with a momentum of 0.9 and weight decay of 1e-4. The CIFAR-10 and CIFAR-100 experiments were each conducted for 200 epochs, with batch sizes of 16 and 200, respectively. For ImageNet, we trained for 400 epochs with a batch size of 16. The initial learning rate was set to 1e-4 for CIFAR-10 and CIFAR-100, and 1e-5 for ImageNet, with multi-step learning rate decay applied. The decay steps were set to 100, 200, and 300 epochs for CIFAR, and at 350 and 400 epochs for ImageNet, with a decay rate of 0.1. We compared our method with existing ZSQ methods [41, 10, 36, 5, 1, 14, 32, 23, 26]. For data generation, we followed the AdaDFQ [36] approach based on ACGAN [35]. Further implementation details are provided in the supplementary material, Section B. In our ablation study, layer-wise quantization was applied to all layers containing activation functions, while channel-wise quantization was performed per channel at the batch level.
CIFAR-100 ResNet-20 (70.33%)
Method GDFQ Qimera+AIT AdaDFQ AdaDFQ+AKT
Baseline +GranQ Baseline +GranQ Baseline +GranQ Baseline +GranQ
3w3a 47.61 59.04+11.43 45.70 60.42+14.72 52.74 62.73+9.99 54.68 62.01+7.33
4w4a 63.39 66.97+3.58 65.80 68.08+2.28 66.81 68.79+1.98 66.94 68.77+1.83
5w5a 66.12 68.96+2.84 69.26 70.14+0.88 69.93 70.05+0.12 69.75 70.21+0.46
[…]
Layer-Q Channel-Q
Layer-Channel
Layer-Channel
(Vectorization, GranQ)
Accuracy (%)
49.98 (±$\pm 0.24$)
51.26 (±$\pm 0.28$)
62.68 (±$\pm 0.11$)
62.73 (±$\pm 0.17$)
Table 5: Comparative analysis of four quantization approaches on ResNet-20 accuracy. The experiments were conducted using the AdaDFQ method on CIFAR-100 with 3-bit quantization. Results are averaged over three repetitions with standard deviation.
5.2 Performance Evaluation
We evaluated the performance of GranQ against SOTA ZSQ methods, with results summarized in Table 3. All comparison experiments were conducted under 3, 4, and 5-bit quantization settings.
CIFAR-10/100. As indicated in Table 3, GranQ consistently achieved the highest accuracy across all bitwidths. For CIFAR-10, it attained accuracies of 94.06% (5-bit), 93.52% (4-bit), and 91.37% (3-bit). For CIFAR-100, the results were 70.05% (5-bit), 68.79% (4-bit), and 62.73% (3-bit). Notably, GranQ outperformed SynQ [26], the previous SOTA, by +5.45% in the CIFAR-100 3-bit setting. This result demonstrates its ability to effectively overcome the limitations of conventional low-bitwidth quantization techniques. Overall, GranQ consistently outperformed existing methods across all bitwidths in both CIFAR-10 and CIFAR-100, with particularly strong improvements in the 3-bit quantization setting. Remarkably, in the CIFAR-10 5-bit setting, GranQ even surpassed the performance of the FP model. These results suggest that we can effectively apply GranQ to small- and medium-scale datasets.
ImageNet. In the ImageNet experiments, GranQ achieved competitive performance across various bitwidths. Specifically, for ResNet-18, it attained top accuracies of 70.39% (4-bit) and 71.31% (5-bit). For ResNet-50, it achieved the highest accuracies of 76.63% (4-bit) and 71.31% (5-bit). Additionally, in the MobileNetV2 setting, GranQ achieved SOTA performance across all bitwidths (3, 4, and 5-bit). In the 3-bit setting, GranQ achieved the second-best accuracies, with 64.41% on ResNet-18 and 70.76% on ResNet-50. These results are -3.77% and -3.23% lower than those of GenQ, respectively. However, GenQ uses a pre-trained diffusion model [22] and prompts during data generation, leading to slower speeds compared with those of typical GAN-based methods [41]. In contrast, GranQ maintains the speed advantage of GAN-based generation while achieving high performance through fine-grained adjustments.
5.3 Ablation Study
5.3.1 Effectiveness Evaluation
GranQ consistently demonstrates performance improvements when applied to various ZSQ methods. As summarized in Table 4, GranQ achieves steady performance improvements when integrated with existing SOTA ZSQ techniques.
First, we analyzed the impact of GranQ on data synthesis-based quantization methods, specifically GDFQ [41] and AdaDFQ [36]. The integration of GranQ into GDFQ [41] led to a notable improvement, with accuracy rising from 47.61% to 59.04% (+11.43%) in the 3-bit setting and from 63.39% to 66.97% (+3.58%) in the 4-bit setting. Similarly, AdaDFQ [36] exhibited improvements of +9.99% in the 3-bit setting (from 52.74% to 62.73%) and +1.98% in the 4-bit setting (from 66.81% to 68.79%). These results suggest that GranQ effectively reduces quantization errors when combined with data synthesis-based quantization methods, leading to enhanced model performance.
Furthermore, GranQ also exhibits significant improvements in methods focused on training quantized models, such as Qimera+AIT [10] and AdaDFQ+AKT [23]. For Qimera+AIT [10], the 3-bit accuracy increased from 45.70% to 60.42% (+14.72%). Similarly, for AdaDFQ+AKT [23], the accuracy improved from 54.68% to 62.01% (+7.33%). These findings demonstrate that GranQ is not only effective in data synthesis-based methods but also enhances performance during the model training process.
Figure 5: Ablation study on the latency of the quantization process in ResNet-20 across different batch sizes. The experiments were conducted using the AdaDFQ method on CIFAR-100 with a 3-bit quantization setting.
5.3.2 Efficiency Evaluation
In general, performing precise quantization that considers both layers and channels leads to an exponential increase in computational cost. However, GranQ is designed to minimize this overhead while maintaining high accuracy by leveraging vectorized operations. To validate this, we measured the latency across various batch sizes (16, 32, 64, 128, and 200). Figure 5 compares four methods.
The first two methods are conventional quantization techniques: layer- and channel-wise. The third is scalar-based iterative, which considers all channels in each layer but performs multiple iterations. The fourth method is GranQ, which integrates both layer and channel information while employing vectorized quantization. Although scalar-based iterative quantization performs the same operations as GranQ, it relies on repeated scalar computations, as illustrated in Equations 1, 2, and 3. In contrast, GranQ enhances efficiency through parallelized computations (Algorithm 1 in supplementary material).
A key observation from Figure 5 is that GranQ induces only a slight increase in latency compared with layer- and channel-wise quantization. Moreover, as summarized in Table 5, GranQ achieves significantly higher accuracy than traditional quantization methods. Even when compared with the less efficient iteration-based layer-channel quantization, GranQ maintains nearly identical accuracy. Additionally, as summarized in Table 2, GranQ substantially reduces quantization error and effectively preserves activations. These findings confirm that GranQ achieves fine-grained quantization while benefiting from parallel computation.
6 Discussion
6.1 Why GranQ is Effective?
The most common activation quantization approaches are known for their low computational cost and fast processing speeds. However, they often fail to accurately capture activation distributions, resulting in substantial quantization errors (Figure 3). In layer-wise quantization, representing the entire activation range with a single scaling factor makes it difficult to reflect fine-grained distribution changes. Similarly, channel-wise quantization encounters challenges owing to significant range differences in activations across layers for the same channel, leading to increased information loss (Section D in the supplementary material). These limitations become more pronounced in low-bit settings, where higher quantization errors result in performance degradation.
To overcome these issues, GranQ introduces the layer-channel scaling method, which combines layer- and channel-wise quantization. Unlike conventional methods, GranQ considers the activation distribution of each channel individually while also incorporating layer information, enabling more precise quantization. Furthermore, although conventional quantization relies on scalar operations with a single scaling factor, GranQ utilizes multiple scaling factors and performs vectorized computations. This approach leverages GPU resources efficiently, minimizing computational overhead while achieving effective quantization.
Experimental results demonstrate that GranQ outperforms traditional layer- and channel-wise quantization methods, achieving higher cosine similarity and lower relative error (Table 2). This demonstrates that GranQ accurately preserves the original activation distribution. Additionally, GranQ maintains computational efficiency in terms of latency and effectively balances the trade-off between computational cost and performance (Figure 5 and Table 5). Most notably, GranQ achieves SOTA performance compared with those of existing layer-wise quantization methods (Table 3), highlighting its efficiency and superiority over scalar iteration approaches.
6.2 Limitation and Future-work
As GranQ aims to precisely preserve activations during quantization, it partially depends on generated data quality. Consequently, studies on data generation remain highly important. The proposed method has the potential to achieve even better performance when applied to high-quality data generation strategies. In future work, we aim to improve the quantization performance by focusing on both the accurate generation of original activations and their effective preservation.
7 Conclusion
We present a comprehensive analysis of activation quantization errors in ZSQ and propose GranQ, a novel quantization method to effectively reduce these errors. Conventional layer- and channel-wise quantization methods rely on a single scaling factor, which often results in substantial activation loss. In contrast, GranQ employs a fine-grained quantization approach that simultaneously considers both layer and channel information, thereby reducing activation loss and enabling more precise quantization. Moreover, GranQ successfully maintains computational efficiency through vectorized quantization while performing detailed and fine-grained quantization.
Through extensive experiments, we demonstrate that GranQ achieves superior performance while preserving computational efficiency. Compared with various SOTA methods, GranQ demonstrates competitive performance and achieves notable improvements. Our study highlights the importance of activation preservation in ZSQ and proposes GranQ as an insightful direction to address this challenge.
Acknowledgement This research was supported by the National Research Foundation (NRF) funded by the Korean government (MSIT) (No. RS-2023-00229822).
Supplementary Material
Appendix A Notation
Symbol Definition
$x$ Original activation value
$x_{q}$ Quantized activation value
$x_{\text{deq}}$ Dequantized activation value
$x_{\min}$, $x_{\max}$ Minimum and maximum activation values
$Q$ Quantization operator
$s$ Scaling factor
$z$ Zero-point for quantization
$b$ Number of quantization bits
$\min\limits_{h,w}(x)$, $\max\limits_{h,w}(x)$ Channel-wise min/max activations
$\mathbb{R}^{C\times H\times W}$ Tensor of shape $C\times H\times W$ in $\mathbb{R}$
[…]
$\|\cdot\|_{2}$ L2 norm (Euclidean norm)
$\in$ Set membership (”element of”)
Mathematical Expressions
Forward Quantization $x_{q}=\left\lfloor\frac{x}{s}+z\right\rceil$
Dequantization $x_{\text{deq}}=s(x_{q}-z)$
Scaling Factor $s=\frac{x_{\max}-x_{\min}}{2^{b}-1}$
Zero-point $z=\left\lfloor\frac{-x_{\min}}{s}\right\rceil$
Cosine Similarity $\frac{X\cdot Q}{\|X\|_{2}\|Q\|_{2}}$
Relative Error $\frac{\|X-Q\|_{2}}{\|X\|_{2}}$
Table 6: Notation for activation quantization equations, algorithm, and evaluation metrics.
In this section, we provide an overview of the main symbols and definitions employed throughout the paper. It specifically includes essential elements such as the original activation ($x$), quantized activation ($x_{q}$), scaling factor ($s$), and zero-point ($z$), along with their corresponding mathematical representations.
Appendix B Details of Experimental Setup
We provide a detailed description of the datasets, experimental setup, and comparison methods utilized in our experiments. This ensures the implementation of the experiments and clearly highlights the strengths of the proposed approach over existing methods.
B.1 Datasets
This paper outlines the experimental settings for CIFAR-10 [9], CIFAR-100 [9], and ImageNet (ILSVRC-2012) [4] datasets. CIFAR-10 and CIFAR-100 contain 10 and 100 classes, respectively, with 32 × 32 resolution color images, and both have 50 K training and 10 K test images. ImageNet comprises 1,000 classes, with 1.2 M training and 50 K validation images, all resized to 224 × 224 for our experiments. We employ these widely used benchmarks for zero-shot quantization (ZSQ) to assess and compare the performance of GranQ against those of existing approaches.
B.2 Experimental Setup
The data generation process followed the settings of AdaDFQ [14]. Specifically, ACGAN [12] was used for data generation with the Adam optimizer, set with a momentum of 0.9 and a learning rate of 1e-3. All hyper-parameters such as $\alpha_{ds}$, $\alpha_{as}$, $\lambda_{l}$, and $\lambda_{u}$ followed the default settings from AdaDFQ [14]. For the fine-tuning process, a warmup phase was applied to ensure training stability, with 4 epochs out of 200 for CIFAR and 50 epochs out of 400 for ImageNet. The batch size was set to 16 for CIFAR-10 and ImageNet, and 200 for CIFAR-100. All experiments used Nesterov-based SGD [11] as the optimizer, with a momentum of 0.9 and a weight decay of 1e-4. In CIFAR experiments, the weight decay of 1e-4 was reduced by a factor of 0.1 every 100 epochs, whereas for ImageNet, decay was applied at 350 and 400 epochs (for ResNet-50 at 150, 300, and 400 epochs). GranQ was implemented using PyTorch [13], and visualizations were generated with matplotlib [7]. All experiments were conducted on an NVIDIA RTX 3090 GPU.
B.3 Comparison Methods
GranQ is a quantization-aware training (QAT) method for zero-shot quantization (ZSQ). Therefore, to evaluate the proposed approach, we conducted comparative experiments using the following state-of-the-art methods:
Algorithm 1 GranQ: Layer-Channel Quantization
1:Activation tensors $\{x_{l}\}_{l=1}^{L}$, where $x_{l}\in\mathbb{R}^{C\times(HW)}$, quantization bit $b$
2:Dequantized activation tensors $\{x_{\text{deq},l}\}_{l=1}^{L}$
3:for each layer $l$ in $L$ do $\triangleright$ Process each layer separately
4:       Compute per-channel min/max values:
$\vec{x}_{\min}=\min_{h,w}(x_{l}),\quad\vec{x}_{\max}=\max_{h,w}(x_{l})$
$\triangleright$ Compute along the channel axis for layer $l$
5:       Compute scaling factors and zero-points (Vectorized):
$\vec{s}=\frac{\vec{x}_{\max}-\vec{x}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{x}_{\min}}{\vec{s}}\right\rceil$
6:       Quantize values (Vectorized):
$x_{q,l}=\left\lfloor\frac{x_{l}}{\vec{s}}+\vec{z}\right\rceil$
$\triangleright$ Element-wise quantization for all channels in layer $l$
7:       Dequantize values (Vectorized):
$x_{\text{deq},l}=\vec{s}(x_{q,l}-\vec{z})$
$\triangleright$ Reconstruct activation tensor for layer $l$
8:end for
9:Return the dequantized tensors:
$\textbf{return}\ \{x_{\text{deq},l}\}_{l=1}^{L}$
  • GDFQ [15]: GDFQ (Generative low-bit data-free quantization) is the first approach to train quantized models using synthetic data through knowledge distillation. It demonstrated that effective quantization is possible even without access to original data.
  • AIT [3]: AIT was the first to show the negative impact of cross-entropy loss when training quantized models with distillation. It also provided a systematic analysis of training strategies in the ZSQ domain.
  • AdaDFQ [14]: AdaDFQ (Adaptive data-free quantization) highlighted the generalization issues of generated data under various quantization settings and proposed a more adaptable data generation method.
  • TexQ [2]: TexQ introduced a data generation method that preserves texture distributions, thereby enhancing the quality of generated data in ZSQ environments.
  • RIS [1]: RIS (Robustness-guided image synthesis) addressed the low semantic quality of generated data and proposed using perturbations to improve data diversity.
  • PLF [5]: PLF was the first to evaluate generated data before quantization, offering a method to ensure more reliable data.
  • GenQ [10]: GenQ introduced the use of stable diffusion in ZSQ, enabling the generation of high-resolution, realistic data.
  • AKT [6]: AKT (Advanced knowledge transfer) proposed refining feature information during distillation to improve the training of quantized models.
  • SynQ [8]: SynQ enhanced the training of quantized models by reducing noise in generated data using low-pass filtering and aligning activation maps.
Figure 6: Ablation study on relative error in ResNet-20 across layers. Layer-Q, Channel-Q, and GranQ represent layer-wise, channel-wise, and proposed quantization. Experiments used AdaDFQ on CIFAR-100 with 3-bit quantization.
Appendix C Algorithm
GranQ is a granular quantization method that processes activation values across all channels within each layer. Algorithm 1 describes the procedure for quantizing the activation tensor $\{x_{deq,l}\}^{L}_{l=1}$. The activation tensor of layer $l$ has a shape of $C\times(HW)$, and the minimum and maximum values are calculated independently for each channel. Based on these values, a scaling factor and zero-point are computed per channel for the quantization process. The resulting quantized values are dequantized back to floating-point representation prior to returning.
Appendix D Extra Analysis
D.1 Analysis of Activation Preservation
To analyze the severe quantization error (activation distortion) occurring in specific layers, we conducted additional experiments, as illustrated in Figure 6. This figure presents the relative error for each layer of ResNet-20 and reveals that certain layers such as the 18th exhibit significantly high errors. These findings suggest that both layer- and channel-wise quantization fail to effectively preserve activations in specific layers. In contrast, GranQ achieves consistently low errors across all layers, ensuring better activation preservation.
D.2 Analysis of Activation Magnitude
In Figure 7, we observe that channel-wise quantization is ineffective in ZSQ settings, similar to layer-wise quantization. As shown in Figure 7(a), the full-precision (FP) model exhibits significant activation magnitude differences across layers, even within the same channel axis. This highlights the limitations of channel-wise quantization in handling quantization errors. Conversely, GranQ addresses this issue by accounting for each channel across all layers, thereby minimizing quantization errors. This is particularly important in the final layers, where maintaining the activations of the FP model is critical for accurate inference. Whereas (a) fails to preserve these distributions, (b) successfully maintains them.
(a) Layer Quantization
(b) GranQ (Ours)
Figure 7: Visualization of activation magnitudes across layers in ResNet-20. (a) illustrates the activation magnitudes with layer-wise quantization, whereas (b) presents the results after applying the proposed GranQ method. In each subplot, the left side depicts activation magnitudes before quantization (32-bit), whereas the right side shows them after quantization (3-bit). The proposed GranQ effectively preserves the activation distribution, mitigating distortions caused by quantization.