### v4

# GranQ: Granular Zero-Shot Quantization with Channel-Wise Activation Scaling in QAT

Abstract
Zero-shot quantization (ZSQ) enables neural network compression without original training data, making it a promising solution for restricted data access scenarios. To compensate for the lack of data, recent ZSQ methods typically rely on synthetic inputs generated from the full-precision model. However, these synthetic inputs often lead to activation distortion, especially under low-bit settings. As a result, existing methods struggle to mitigate this issue due to coarse activation scaling. To address this issue, we propose GranQ, a novel activation quantization framework that efficiently applies per-channel scaling through vectorized computation. In contrast to conventional channel-wise methods, which apply vectorization only to the quantization step, GranQ improves efficiency by vectorizing the scaling operation. This design allows GranQ to maintain fine-grained quantization granularity with minimal computational overhead, even in low-bit environments. Extensive experiments under quantization-aware training (QAT) settings demonstrate that GranQ consistently outperforms state-of-the-art ZSQ methods across CIFAR and ImageNet. In particular, our method achieves up to 5.45% higher accuracy in the 3-bit setting on CIFAR-100 and even surpasses the full-precision baseline on CIFAR-10. Furthermore, GranQ achieves significant speedup in quantization latency over conventional per-channel methods, demonstrating improved efficiency. With these findings, we anticipate that GranQ will inspire future research beyond conventional ZSQ approaches centered on data generation and model fine-tuning.
1 Introduction
Neural network compression has been extensively studied for the practical deployment of large-scale deep learning (DL) models. In particular, reducing model size while minimizing performance degradation is crucial for utilizing DL on edge devices (e.g., mobile phones, embedded systems, and drones). Major approaches to model compression include quantization [1, 2, 3], pruning [4, 5, 6, 7], knowledge distillation [8, 9, 10], and neural architecture search [11, 12], as surveyed in [13, 14]. Among these, quantization has emerged as the most actively studied technique. It serves as an effective compression method by reducing unnecessary representational ranges in the model. However, it often requires fine-tuning or calibration to match full-precision (FP) model performance [3]. To address this, zero-shot quantization (ZSQ), also known as data-free quantization, has been proposed to compress models without the original training data. Since the introduction of ZeroQ [15], studies on ZSQ have advanced in two main directions. The first direction focuses on data generation, where synthetic data are created from the FP model. The second direction focuses on effectively applying the activation distributions of the synthetic data to the quantized (Q) model. This second direction is further divided into post-training quantization (PTQ), which calibrates the activation distributions, and quantization-aware training (QAT), which fine-tunes the Q model directly. We categorize existing ZSQ methods based on these directions, as summarized in Table 1.
First, data generation studies focus on generating high-quality data to effectively train the Q model [16, 17, 18]. Meanwhile, calibration (PTQ) studies aim to minimize the quantization error by calibrating the Q model without additional training [19, 20]. Finally, fine-tuning (QAT) studies focus on transferring key information from the FP model to the Q model through knowledge distillation [21, 22, 23, 24].
However, despite extensive studies, severe performance degradation in low-bit quantization remains unresolved. To address this issue, we performed an in-depth analysis of the ZSQ process, focusing on why low-bitwidth settings still suffer from performance loss even after QAT fine-tuning. Our findings reveal that quantization errors mainly stem from the loss of activation values instead of data quality or training methods. Notably, we found that layer-wise (per-tensor) quantization is no longer suitable for preserving activations in ZSQ, as it leads to coarse and inaccurate representations.
Based on this analysis, we introduce GranQ, a novel ZSQ approach that enables channel-wise activation quantization by adapting the scale to each channel of every input activation. This allows us to overcome the limitations of conventional activation quantization. This dynamic adjustment minimizes activation loss and preserves the original activation values by reducing quantization errors, as shown in Figure 1. The proposed method effectively handles activation loss in low-bit quantization and achieves state-of-the-art (SOTA) performance in QAT settings on the CIFAR and ImageNet datasets. Furthermore, We apply vectorization to the scaling operation, which is not applied in conventional channel-wise quantization, reducing latency while enabling fine-grained activation quantization. Our contributions can be summarized as follows:
(a) Layer-wise Quantization
(b) GranQ (Channel-wise, Ours)
Figure 1: Comparison between (a) layer-wise quantization and (b) GranQ on the CIFAR-10. Each subfigure visualizes the 32-bit FP (left) and 3-bit quantized (right) activations of the first ResNet-20 layer. GranQ better preserves the original activation with minimal distortion.
  • We identify critical limitations of layer-wise activation quantization in low-bit ZSQ. Our findings reveal that conventional activation quantization methods relying on layer-wise granularity suffer from significant activation loss. This limitation becomes more severe in ZSQ settings with synthetic data.
  • We propose GranQ, a novel method that supports granular quantization and maintains computational efficiency. We introduce a fine-grained activation quantization mechanism that considers every channel to reduce activation loss. Although per-channel activation quantization is known to improve precision, it has not been widely adopted due to its high computational cost. We address this by applying vectorization to the scaling operation, enabling efficient per-channel quantization with minimal overhead. To the best of our knowledge, our approach is the first to address the ZSQ problem.
  • We achieve SOTA performance over existing ZSQ methods through extensive evaluation. Specifically, on the CIFAR-100 dataset, in the 3-bit quantization setting, GranQ achieves an accuracy of 62.73%, improving by 5.45% over the latest method on ResNet-20. Furthermore, on the CIFAR-10 dataset in the 5-bit quantization setting, GranQ achieves an accuracy of 94.06%, slightly exceeding the FP model performance by 0.17%.
2 Related Work
2.1 Quantization
Quantization reduces the representational range of deep neural networks (DNNs) and minimizes memory usage. It is typically categorized into post-training quantization (PTQ) and quantization-aware training (QAT) [3, 25]. PTQ applies quantization after training without further updates, which makes it efficient and easy to implement. However, it is sensitive to scaling errors and often relies on calibration to improve accuracy. In contrast, QAT incorporates quantization during training and optimizes quantized activations using the straight-through estimator [3, 26]. Both methods require data—PTQ for calibration and QAT for fine-tuning. In practice, access to training data is often limited. To address this, zero-shot quantization (ZSQ) has been proposed to perform quantization without using original data [16].
Data Generation
(PTQ, QAT)
Calibration
Fine-tuning
ZeroQ (CVPR 20) [15]
GDFQ (ECCV 20) [16]
DSG (CVPR 21) [27]
MixMix (CVPR 21) [28]
Qimera (NeurIPS 21) [29]
IntraQ (CVPR 22) [30]
GENIE (CVPR 23) [31]
AdaDFQ (CVPR 23) [17]
Casual-DFQ (ICCV 23) [32]
TexQ (NeurIPS 23) [33]
RIS (AAAI 24) [34]
GenQ (ECCV 24) [18]
SQuant (ICLR 22) [19]
UDFC (ICCV 23) [20]
AIT (CVPR 22) [21]
PLF (CVPR 24) [22]
SynQ (ICLR 25) [24]
GranQ (Ours)
Table 1: Categorizing zero-shot quantization algorithms. PTQ denotes post-training quantization, and QAT denotes quantization-aware training. Data generation methods typically support both PTQ and QAT, whereas violet-marked methods utilize fine-tuning for data generation and support only QAT.
2.2 Zero-shot Quantization
As summarized in Table 1, ZSQ was initially introduced by ZeroQ [15], leading to the development of various ZSQ algorithms. These studies have explored diverse approaches to improve quantization performance under data-free settings. Recent studies have focused on generating high-quality synthetic data [31, 18] and developing more effective methods to utilize them [21, 23, 24]. However, current ZSQ methods heavily rely on augmented synthetic inputs. This causes large variation in activation scales across channels. In practice, such input-dependent channel-wise quantization is rarely adopted due to its high computational cost. Therefore, recent studies have applied channel-wise activation quantization either in a limited conditions within QAT settings [35], or exclusively in PTQ scenarios with access to a small amount of data [36, 37]. This highlights the need for efficient channel-wise activation quantization for QAT in ZSQ.
3 Preliminaries and Problem Definition
In this section, we provide a new definition of the existing ZSQ problem and introduce the preliminaries. The detailed mathematical notations are listed in Table 6 in the appendix.
3.1 Activation Quantization
Activation quantization reduces the precision of intermediate activation values by converting them into low-bitwidth integers [38]. It is commonly used with weight quantization to compress models, and the linear quantization scheme is most widely adopted. This process typically consists of quantization, scaling parameter computation, and dequantization [1, 3, 39].
The quantization operator $\mathcal{Q}$ maps a floating-point value $x$ into an integer $x_{q}$ using a scaling factor $s$ and zero-point $z$:
$x_{q}=\mathcal{Q}(x,s,z)=\left\lfloor\frac{x}{s}+z\right\rceil$ (1)
Here, $s$ normalizes the activation range, and $z$ shifts the scaled value to account for asymmetric quantization. These parameters are computed as:
$s=\frac{x_{\max}-x_{\min}}{2^{b}-1},\quad z=\left\lfloor\frac{-x_{\min}}{s}\right\rceil$ (2)
where $x_{\min}$ and $x_{\max}$ denote the range of activation values, and $b$ is the bit-width. A large range yields a large $s$, increasing quantization error; a smaller range allows finer resolution.
During inference, the dequantization step restores quantized values to real numbers:
$x_{\text{deq}}=s(x_{q}-z)$ (3)
This enables recovery of activation information needed for final predictions.
3.2 Problem Definition
Existing activation quantization methods are designed based on layer-wise quantization to minimize computational cost. However, a fixed scaling factor struggles to handle varying activations. This issue becomes more severe under low-bitwidth settings in ZSQ environments. The following outlines the primary challenges faced in ZSQ under low-bitwidth settings.
(Problem 1) Coarse Activation Quantization by Single-Range Scaling
Conventional activation quantization uses a single scaling range per layer for efficiency, which was acceptable at higher bitwidths or when real data is available. However, this design becomes problematic in ZSQ, where range adjustment relies on synthetic data. These synthetic distributions are often biased or irregular, making precise quantization more difficult.
(Problem 2) Severe Quantization Error in Certain Layers
During activation quantization, certain layers induce severe quantization errors, leading to significant activation distortions. These distortions substantially differ from the original distribution and can directly impact the performance of the model.
4 Observation and Methodology
We observe the causes of activation distortion during quantization and analyze how to mitigate them. Based on this, we introduce the proposed method known as GranQ.
Figure 2: Activation magnitudes of ResNet-20 on CIFAR-10 using 32-bit FP (left) and 3-bit layer-wise quantization (right). Severe distortions appear, especially in the final layer.
4.1 Observation
Layer-wise quantization effectively captures the activation distribution of each layer. However, it becomes inefficient when activation distributions within a layer exhibit significant variation along the channel axis. Particularly in ZSQ, as shown in the FP graph in Figure 1(a), the activation distributions can differ substantially across channels even within the same layer. This implies that limited representation bits are forced to capture a wide range of activation values, making fine-grained quantization difficult. It consistently appears across all layers. We analyzed this phenomenon occurring across all layers and provided a summary in Table 2. It presents the cosine similarity $\frac{X\cdot Q}{|X|_{2}|Q|_{2}}$, along with the relative error $\frac{|X-Q|_{2}}{|X|_{2}}$. From the results, we observe that layer-wise quantization fails to preserve activation information effectively. This reveals that using a single scaling range for activation quantization has a critical limitation, particularly in ZSQ. In contrast, GranQ, which applies scaling per channel, achieves 1.34$\times$ higher similarity and 2.94$\times$ lower relative error than layer-wise quantization, demonstrating its effectiveness in preserving activation information.
Furthermore, Figure 2 provides a visual representation of the effect of quantization on activation values across layers. In layer-wise quantizaiton, although the overall activations generally reflect the FP activations to some extent, certain layers exhibit fine activation values being mapped to discrete values, leading to coarse representations. These coarse representations have been identified as a major factor leading to a decline in the overall expressiveness and inference performance of the model. Further analysis of Figure 2 is provided in Figure 5 and in Appendix (Figure 7).
Activaiton Quantization Layer-wise Channel-wise (GranQ)
$Avg.\frac{X\cdot Q}{\|X\|_{2}\|Q\|_{2}}(\uparrow)$
$Avg.\frac{\|X-Q\|_{2}}{\|X\|_{2}}(\downarrow)$
0.5111
0.3129
0.6835
0.1063
Table 2: Average cosine similarity and relative error for 3-bit activation quantization on ResNet-20 (CIFAR-100). GranQ (channel-wise) outperforms layer-wise quantization by achieving less distortion in activation representation, as indicated by higher similarity and lower error.
4.2 Methodology
Based on our analysis, we propose GranQ, a fine-grained ZSQ method designed to reduce computational overhead (Figure 3). GranQ applies channel-wise scaling through activation reshaping and decomposition, and achieves fast parallel processing via vectorization of scaling parameters, which enables efficient computation of both scaling and quantization.
4.2.1 Activation Reshaping and Decomposition
We propose activation reshaping and decomposition, which reshapes each activation map from a three-dimensional tensor of shape $(C\times H\times W)$ into a two-dimensional matrix of shape $(C\times HW)$. This transformation enables decomposition of the activation map along the channel axis, allowing each channel to be processed independently. It also facilitates vectorized computation of per-channel statistics (e.g., min and max) required for scaling factor calculation during quantization. The activation reshaping and decomposition is defined as follows:
$\begin{gathered}A_{l}=\left[A_{l}(1,:,:),\,A_{l}(2,:,:),\,\dots,\,A_{l}(C,:,:)\right]\in\mathbb{R}^{C\times H\times W}\\ \vec{A}_{\min}=\left[\min_{(h,w)}A_{l}(c,h,w)\right]_{c=1}^{C}\in\mathbb{R}^{C}\\ \vec{A}_{\max}=\left[\max_{(h,w)}A_{l}(c,h,w)\right]_{c=1}^{C}\in\mathbb{R}^{C}\end{gathered}$ (4)
[…]
4.2.2 Vectorization of Scaling Parameters
In the vectorization of scaling parameters stage, multiple scaling ranges are stored in a vectorized form for each channel. This allows dynamic adaptation to the diverse activation distributions induced by synthetic inputs. Thus, GranQ contrasts with traditional per-channel quantization, which uses fixed scaling parameters for each channel regardless of input variation. As shown in Equation 5, the scaling factors $\vec{s}$ and zero-points $\vec{z}$ are computed in vectorized form. The activation values are then quantized and dequantized according to Equation 6, with all operations applied in parallel across channels.
$\vec{s}=\frac{\vec{A}_{\max}-\vec{A}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{A}_{\min}}{\vec{s}}\right\rceil$ (5)
$A_{q,l}=\left\lfloor\frac{A}{\vec{s}}+\vec{z}\right\rceil,A_{\text{deq},l}=\vec{s}(A_{q,l}-\vec{z})$ (6)
This process allows efficient and adaptive quantization of activation tensors exhibiting high variation across inputs, leading to improved computational efficiency (Table 7 in appendix).
[…]
5 Experiments
In this section, we thoroughly evaluate the effectiveness of GranQ. Experiments are conducted on diverse benchmark datasets, with the performance compared with those of existing ZSQ methods.
5.1 Experimental Setup and Details
The experiments were conducted using widely adopted ZSQ evaluation datasets, including CIFAR-10, CIFAR-100 [40], and ImageNet (ILSVRC 2012) [41] validation datasets. For the CIFAR datasets, ResNet-20 [42] was used as the quantization model, whereas ResNet-18 [42], ResNet-50 [42], and MobileNetV2 [43] were employed for ImageNet. All experiments were conducted using the SGD optimizer [44] with a momentum of 0.9 and weight decay of 1e-4. The CIFAR-10 and CIFAR-100 experiments were each conducted for 200 epochs, with batch sizes of 16 and 200, respectively. For ImageNet, we trained for 400 epochs with a batch size of 16. The initial learning rate was set to 1e-4 for CIFAR-10 and CIFAR-100, and 1e-5 for ImageNet, with multi-step learning rate decay applied. The decay steps were set to 100, 200, and 300 epochs for CIFAR, and at 350 and 400 epochs for ImageNet, with a decay rate of 0.1. We compared our method with existing ZSQ methods [16, 21, 17, 33, 34, 22, 18, 23, 24]. For data generation, we followed the AdaDFQ [17] approach based on ACGAN [45]. Further implementation details are provided in the appendix, Section A.3.2 and A.3.3. In our ablation study, layer-wise quantization was applied to all layers containing activation functions, while channel-wise quantization was performed per channel at the batch level.
Dataset
Model
(FP 32)
Bits
GDFQ
(ECCV 20)
ARC+AIT
(CVPR 22)
AdaDFQ
(CVPR 23)
TexQ
(NeurIPS 24)
AIT+RIS
(AAAI 24)
GenQ
(ECCV 24)
AKT
(SAC 25)
SynQ
(ICLR 25)
GranQ
[…]
Table 3: Accuracy Evaluation of QAT Methods for ZSQ. $w$ and $a$ represent weight and activation, respectively. Bold values indicate the best accuracy, and underlined values denote the second-best accuracy. $\dagger$ indicates our re-implementation with the official code.
CIFAR-100 ResNet-20 (70.33%)
Method GDFQ Qimera+AIT AdaDFQ AdaDFQ+AKT
Baseline +GranQ Baseline +GranQ Baseline +GranQ Baseline +GranQ
[…]
Table 4: Accuracy of existing SOTA methods with the integration of GranQ. GDFQ and AdaDFQ focus on data generation, whereas Qimera+AIT and AdaDFQ+AKT primarily enhance quantized model training.
Figure 4: Latency of ResNet-20 quantization across batch sizes on CIFAR-100 with 3-bit setting.
Figure 5: Relative quantization error across layers in ResNet-20 with 3-bit quantization on CIFAR-100.
5.2 Performance Evaluation
We evaluated the performance of GranQ against SOTA ZSQ methods, with results summarized in Table 3. All comparison experiments were conducted under 3, 4, and 5-bit quantization settings.
CIFAR-10/100. As indicated in Table 3, GranQ consistently achieved the highest accuracy across all bitwidths. For CIFAR-10, it attained accuracies of 94.06% (5-bit), 93.52% (4-bit), and 91.37% (3-bit). For CIFAR-100, the results were 70.05% (5-bit), 68.79% (4-bit), and 62.73% (3-bit). Notably, GranQ outperformed SynQ [24], the previous SOTA, by +5.45% in the CIFAR-100 3-bit setting. This result demonstrates its ability to effectively overcome the limitations of conventional low-bitwidth quantization techniques. Overall, GranQ consistently outperformed existing methods across all bitwidths in both CIFAR-10 and CIFAR-100, with particularly strong improvements in the 3-bit quantization setting. Remarkably, in the CIFAR-10 5-bit setting, GranQ even surpassed the performance of the FP model. These results suggest that we can effectively apply GranQ to small- and medium-scale datasets.
ImageNet. In the ImageNet experiments, GranQ achieved competitive performance across various bitwidths. Specifically, for ResNet-18, it attained top accuracies of 70.39% (4-bit) and 71.31% (5-bit). For ResNet-50, it achieved the highest accuracies of 76.63% (4-bit) and 71.31% (5-bit). Additionally, in the MobileNetV2 setting, GranQ achieved SOTA performance across all bitwidths (3, 4, and 5-bit). In the 3-bit setting, GranQ achieved the second-best accuracies, with 64.41% on ResNet-18 and 70.76% on ResNet-50. These results are -3.77% and -3.23% lower than those of GenQ, respectively. However, GenQ uses a pre-trained diffusion model [46] and prompts during data generation, leading to slower speeds compared with those of typical GAN-based methods [16]. In contrast, GranQ improves the quantization mechanism without relying heavily on diffusion-based data generation, as GenQ does. As our method is adaptable to existing QAT frameworks, it can potentially achieve even higher performance when combined with GenQ’s data generation.
5.3 Ablation Study
5.3.1 Effectiveness Evaluation
GranQ consistently demonstrates performance improvements when applied to various ZSQ methods. As summarized in Table 4, GranQ achieves steady performance improvements when integrated with existing SOTA ZSQ techniques.
First, we analyzed the impact of GranQ on data synthesis-based quantization methods, specifically GDFQ [16] and AdaDFQ [17]. The integration of GranQ into GDFQ [16] led to a notable improvement, with accuracy rising from 47.61% to 59.04% (+11.43%) in the 3-bit setting and from 63.39% to 66.97% (+3.58%) in the 4-bit setting. Similarly, AdaDFQ [17] exhibited improvements of +9.99% in the 3-bit setting (from 52.74% to 62.73%) and +1.98% in the 4-bit setting (from 66.81% to 68.79%). These results suggest that GranQ effectively reduces quantization errors when combined with data synthesis-based quantization methods, leading to enhanced model performance.
Furthermore, GranQ also exhibits significant improvements in methods focused on training quantized models, such as Qimera+AIT [21] and AdaDFQ+AKT [23]. For Qimera+AIT [21], the 3-bit accuracy increased from 45.70% to 60.42% (+14.72%). Similarly, for AdaDFQ+AKT [23], the accuracy improved from 54.68% to 62.01% (+7.33%). These findings demonstrate that GranQ is not only effective in data synthesis-based methods but also enhances performance during the model training process.
5.3.2 Efficiency Evaluation
Method
Quantization
Application
Scaling Factor
Computation
Accuracy (%)
Quantization
Latency (ms)
Layer-wise 49.98 ($\pm$ 0.24) 0.227
Channel-wise (Conventional scalar-loop) 62.68 ($\pm$ 0.11) 103.671
Channel-wise (Vectorized, GranQ) 62.73 ($\pm$ 0.17) 0.236
Table 5: Ablation study of vectorization in scaling computation and quantization. GranQ applies vectorization to both stages, achieving high accuracy with low latency. Latency is measured with batch size 16. Experiments use AdaDFQ on ResNet-20 (CIFAR-100, 3-bit).
While fine-grained activation quantization often incurs high computational cost due to per-channel scaling, GranQ maintains efficiency by vectorizing the scaling computation. By reshaping activation tensors, our method eliminates the overhead of iterative channel-wise operations. To assess this in practice, we measured quantization latency across various batch sizes (16, 32, 64, 128, and 200). Figure 5 compares three methods: layer-wise quantization, channel-wise quantization with scalar-loop scaling (Algorithm 2 in appendix), and our method (GranQ), which efficiently parallelizes scaling across channels (Algorithm 1 in appendix).
The two channel-wise methods perform per-channel quantization. The conventional method computes scaling factors sequentially, whereas GranQ vectorizes the scaling computation across channels for improved efficiency. A key observation from Figure 5 is that GranQ incurs only a slight increase in latency compared to conventional layer-wise quantization. With only a minor increase in latency, GranQ achieves higher accuracy than the layer-wise baseline, as shown in Table 5. Compared to scalar-loop-based channel quantization, which implements the same computation sequentially, GranQ achieves nearly identical accuracy. Furthermore, Table 2 shows that GranQ substantially reduces quantization error and more effectively preserves activation information. These results demonstrate that GranQ enables fine-grained quantization with significantly improved efficiency. Additional analysis and experimental results on inference efficiency are provided in Table 7 of the appendix.
6 Discussion
6.1 Why GranQ is Effective?
The most common approaches to activation quantization (i.e., layer-wise quantization) are known for their low computational cost and fast processing speed. However, they often fail to accurately capture activation distributions, resulting in substantial quantization errors (Figure 2). In layer-wise quantization, representing the entire activation range with a single scaling factor makes it difficult to reflect fine-grained distribution changes. These limitations become more severe in low-bit settings, especially when synthetic data exhibit distributional shifts.
To overcome these issues, GranQ computes and stores per-channel scaling factors for each input in a unified vector form, enabling parallel processing across all channels. Unlike conventional methods, it performs quantization based on each input’s channel-wise distribution, resulting in more precise and efficient computation.
6.2 Limitation and Future-work
In ZSQ settings, synthetic data often exhibit broader and more distorted activation distributions compared to original data (Figure 6 in appendix). GranQ addresses this challenge by applying dynamic quantization for each channel based on its activation distribution, making it particularly effective in QAT scenarios. However, GranQ is less suitable for PTQ, which relies on static activation quantization. As future work, we plan to develop quantization methods that preserve the fine-grained scaling of GranQ while enhancing adaptability to PTQ settings.
7 Conclusion
Layer-wise and channel-wise quantization are both commonly used in weight quantization, as weights remain fixed after training. In contrast, channel-wise scaling in activation quantization is more challenging, as activation values vary with each input, making it difficult to apply fine-grained scaling. These limitations become more pronounced in QAT settings, where large amounts of data are used. Under these conditions, quantization is strongly affected by activation distributions. This issue is especially critical in ZSQ, where training data are synthesized using internal statistics from the FP model. Synthetic data generated in this way often produce distorted activation distributions compared to original data, resulting in increased quantization error during scaling. To address these issues, we propose GranQ, a dynamic channel-wise quantization method that adapts scaling factors to each input’s activation distribution. This enables more precise quantization even under distorted distributions, as often observed in ZSQ. In addition, GranQ improves computational efficiency by applying per-input, per-channel scaling through vectorized operations, which is rarely adopted in activation quantization due to its implementation complexity.
Through extensive experiments, we demonstrate that GranQ achieves superior performance while preserving computational efficiency. Compared with various SOTA methods, GranQ demonstrates competitive performance and achieves notable improvements. Our study highlights the importance of activation preservation in ZSQ and proposes GranQ as an insightful direction to address this challenge.
Appendix A Appendix
A.1 Notation
We provide an overview of the main symbols and definitions employed throughout the paper. It specifically includes essential elements such as the original activation ($A$), quantized activation ($A_{q}$), scaling factor ($s$), and zero-point ($z$), along with their corresponding mathematical representations.
Symbol Definition
$A$ Original activation value
$A_{q}$ Quantized activation value
$A_{\text{deq}}$ Dequantized activation value
$A_{\min}$, $A_{\max}$ Minimum and maximum activation values
$Q$ Quantization operator
$s$ Scaling factor
$z$ Zero-point for quantization
$b$ Number of quantization bits
$\min\limits_{h,w}(A)$, $\max\limits_{h,w}(A)$ Channel-wise min/max activations
$\mathbb{R}^{C\times H\times W}$ Tensor of shape $C\times H\times W$ in $\mathbb{R}$
[…]
$\|\cdot\|_{2}$ L2 norm (Euclidean norm)
$\in$ Set membership ("element of")
Mathematical Expressions
Forward Quantization $x_{q}=\left\lfloor\frac{x}{s}+z\right\rceil$
Dequantization $x_{\text{deq}}=s(x_{q}-z)$
Scaling Factor $s=\frac{x_{\max}-x_{\min}}{2^{b}-1}$
Zero-point $z=\left\lfloor\frac{-x_{\min}}{s}\right\rceil$
Cosine Similarity $\frac{X\cdot Q}{\|X\|_{2}\|Q\|_{2}}$
Relative Error $\frac{\|X-Q\|_{2}}{\|X\|_{2}}$
Table 6: Notation for activation quantization equations, algorithm, and evaluation metrics.
A.2 Algorithm
While prior works apply vectorized computation during the quantization step, they typically compute per-channel scaling parameters using iterative channel-wise operations. GranQ differs by also vectorizing the scaling computation, allowing both steps to be processed in parallel. Algorithm 1 outlines the complete process of quantizing and dequantizing activation tensors using GranQ.
In Algorithm 1, both the scaling and quantization steps are implemented using vectorized tensor operations across all channels. Specifically, the per-channel minimum and maximum values are computed simultaneously, and the resulting scaling factors and zero-points are applied through broadcasting. This eliminates the need for channel-wise loops and allows efficient execution on modern hardware accelerators.
In contrast, Algorithm 2 follows a conventional channel-wise quantization approach, where the scaling parameters are computed sequentially for each channel using an explicit for-loop. The quantization and dequantization steps are still vectorized per channel. However, the scaling computation remains scalar and iterative. This leads to increased computational overhead, especially when applied to deep networks with many channels. By removing this bottleneck, GranQ achieves both higher parallelism and lower latency, which are particularly beneficial in low-bitwidth quantization scenarios.
Algorithm 1 GranQ: Vectorized Per-Channel Scaling for Activation Quantization
1:Activation tensors $\{A_{l}\}_{l=1}^{L}$, where $A_{l}\in\mathbb{R}^{C\times H\times W}$, quantization bit $b$
2:Dequantized activation tensors $\{A_{\text{deq},l}\}_{l=1}^{L}$
3:for each layer $l$ in $L$ do
4:    Reshape:
$A_{l}\leftarrow\text{reshape}(A_{l})\in\mathbb{R}^{C\times(HW)}$
$\triangleright$ Flatten spatial dimensions
5:    Decompose by channel:
$\vec{A}_{\min}=\min_{h,w}(A_{l}),\quad\vec{A}_{\max}=\max_{h,w}(A_{l})$
$\triangleright$ Extract per-channel range
6:    Compute scaling parameters:
$\vec{s}=\frac{\vec{A}_{\max}-\vec{A}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{A}_{\min}}{\vec{s}}\right\rceil$
7:    Quantize (vectorized):
$A_{q,l}=\left\lfloor\frac{A_{l}}{\vec{s}}+\vec{z}\right\rceil$
8:    Dequantize (vectorized):
$A_{\text{deq},l}=\vec{s}(A_{q,l}-\vec{z})$
9:end for
10:Return: $\{A_{\text{deq},l}\}_{l=1}^{L}$
Algorithm 2 Conventional Channel-Wise Activation Quantization (Scalar-loop for scaling)
1:Activation tensors $\{A_{l}\}_{l=1}^{L}$, where $A_{l}\in\mathbb{R}^{C\times H\times W}$, quantization bit $b$
2:Dequantized activation tensors $\{A_{\text{deq},l}\}_{l=1}^{L}$
3:for each layer $l$ in $L$ do
4:    Compute per-channel scaling parameters:
5:    for each channel $c$ in $C$ do
$s_{c}=\frac{\max(A_{l,c})-\min(A_{l,c})}{2^{b}-1},\quad z_{c}=\left\lfloor-\frac{\min(A_{l,c})}{s_{c}}\right\rceil$
6:    end for
7:    Apply quantization (vectorized):
$A_{q,l}[c,:,:]=\left\lfloor\frac{A_{l}[c,:,:]}{s_{c}}+z_{c}\right\rceil,\quad\forall c$
8:    Dequantize (vectorized):
$A_{\text{deq},l}[c,:,:]=s_{c}(A_{q,l}[c,:,:]-z_{c}),\quad\forall c$
9:end for
10:Return: $\{A_{\text{deq},l}\}_{l=1}^{L}$
A.3 Details of Experimental Setup
We describe the datasets, experimental setup, and baselines to ensure reproducibility and highlight both the effectiveness and implementation accessibility of our approach.
A.3.1 Datasets
This paper outlines the experimental settings for CIFAR-10 [13], CIFAR-100 [13], and ImageNet (ILSVRC-2012) [14] datasets. CIFAR-10 and CIFAR-100 contain 10 and 100 classes, respectively, with 32 × 32 resolution color images, and both have 50 K training and 10 K test images. ImageNet comprises 1,000 classes, with 1.2 M training and 50 K validation images, all resized to 224 × 224 for our experiments. We employ these widely used benchmarks for zero-shot quantization (ZSQ) to assess and compare the performance of GranQ against those of existing approaches.
A.3.2 Experimental Setup
The data generation process followed the settings of AdaDFQ [1]. Specifically, ACGAN [2] was used for data generation with the Adam optimizer, set with a momentum of 0.9 and a learning rate of 1e-3. All hyper-parameters such as $\alpha_{ds}$, $\alpha_{as}$, $\lambda_{l}$, and $\lambda_{u}$ followed the default settings from AdaDFQ [1]. For the fine-tuning process, a warmup phase was applied to ensure training stability, with 4 epochs out of 200 for CIFAR and 50 epochs out of 400 for ImageNet. The batch size was set to 16 for CIFAR-10 and ImageNet, and 200 for CIFAR-100. All experiments used Nesterov-based SGD [47] as the optimizer, with momentum 0.9 and weight decay 1e-4. In CIFAR experiments, the weight decay of 1e-4 was reduced by a factor of 0.1 every 100 epochs, whereas for ImageNet, decay was applied at 350 and 400 epochs (for ResNet-50 at 150, 300, and 400 epochs). GranQ was implemented using PyTorch [3], and visualizations were generated with matplotlib [4]. All experiments were conducted on an NVIDIA RTX 3090 GPU.
A.3.3 Comparison Methods
GranQ is a quantization-aware training (QAT) method for zero-shot quantization (ZSQ). Therefore, to evaluate the proposed approach, we conducted comparative experiments using the following state-of-the-art methods:
  • GDFQ [5]: GDFQ (Generative low-bit data-free quantization) is the first approach to train quantized models using synthetic data through knowledge distillation. It demonstrated that effective quantization is possible even without access to original data.
  • AIT [6]: AIT was the first to show the negative impact of cross-entropy loss when training quantized models with distillation. It also provided a systematic analysis of training strategies in the ZSQ domain.
  • AdaDFQ [1]: AdaDFQ (Adaptive data-free quantization) highlighted the generalization issues of generated data under various quantization settings and proposed a more adaptable data generation method.
  • TexQ [7]: TexQ introduced a data generation method that preserves texture distributions, thereby enhancing the quality of generated data in ZSQ environments.
  • RIS [8]: RIS (Robustness-guided image synthesis) addressed the low semantic quality of generated data and proposed using perturbations to improve data diversity.
  • GenQ [10]: GenQ introduced the use of stable diffusion in ZSQ, enabling the generation of high-resolution, realistic data.
  • AKT [11]: AKT (Advanced knowledge transfer) proposed refining feature information during distillation to improve the training of quantized models.
  • SynQ [12]: SynQ enhanced the training of quantized models by reducing noise in generated data using low-pass filtering and aligning activation maps.
These methods cover various aspects of ZSQ, including data generation, training objectives. We compare them to demonstrate the effectiveness and generality of GranQ, which focuses on improving the quantization process itself by introducing an optimized, vectorized scaling mechanism. This structural enhancement allows GranQ to be seamlessly integrated into existing QAT-based ZSQ pipelines, providing consistent gains across different settings. Furthermore, as several methods have not yet released official code, we expect that GranQ can be readily applied as their implementations become publicly available.
A.4 Extra Analysis
A.4.1 Data Distribution Analysis
(a) Original inputs (Skewness: 0.2599)
(b) Synthetic inputs (Skewness: 0.4877)
Figure 6: Pixel distribution comparison between (a) original and (b) synthetic inputs on ImageNet. The synthetic inputs, generated by AdaDFQ [17], exhibit higher skewness and a rightward shift, indicating the presence of more extreme activation values and motivating the need for granular scaling in quantization.
Figure 6 compares the pixel distributions of original and synthetic inputs derived from ImageNet. The skewness of the original input is 0.2599, whereas that of the synthetic input increases to 0.4877, with the distribution shifted toward higher pixel values. This indicates that synthetic data, often generated without real image priors, tends to contain a greater number of extreme activation values. As a result, the likelihood of outliers increases, which fixed-range quantization cannot effectively handle. In particular, when a single global range is applied across all channels, the presence of outliers can compress the majority of values into a narrow dynamic range, leading to severe quantization distortion. These observations underscore the limitations of conventional scaling strategies. To address this, we propose a granular, channel-wise scaling approach that dynamically adapts to per-channel statistics and better preserves activation fidelity under distributional mismatch.
A.4.2 Computational Efficiency Analysis
Although GranQ and conventional channel-wise quantization share a higher theoretical complexity of $\mathcal{O}(L\cdot C\cdot H\cdot W)$ compared to the layer-wise baseline with $\mathcal{O}(L\cdot H\cdot W)$, all three methods apply their respective scaling strategies only during the training phase. Since QAT fixes the scaling factors as static parameters prior to deployment, these operations are not executed during inference, resulting in no runtime overhead. Consequently, the per-channel granularity used in GranQ does not incur any additional inference latency. As shown in Table 7, all methods achieve comparable inference latency regardless of the scaling strategy, demonstrating that GranQ maintains runtime efficiency while benefiting from fine-grained scaling during training.
Table 7: Comparison of activation quantization latency on ResNet-18 (ImageNet, batch size = 128) under a 3-bit quantization setting. All latency values are measured per batch and averaged over 3 evaluations, with standard deviation reported in parentheses. Theoretical complexity reflects quantization-time operations and does not affect inference latency, which is measured separately.
Method Theoretical Complexity Scaling Strategy Inference Latency (ms)
Activation (Layer-wise) $O(L\cdot H\cdot W)$ Layer-wise uniform scaling 44.194 (+0.190)
Activation (Channel-wise, Scalar-loop) $O(L\cdot C\cdot H\cdot W)$ Channel-wise scalar scaling (loop-based) 44.272 (+0.158)
GranQ (Channel-wise, Vectorization) $O(L\cdot C\cdot H\cdot W)$ Channel-wise vectorized scaling 44.231 (+0.186)
A.4.3 Analysis of Activation Preservation
We analyzed how activation distortion propagates across layers in a ZSQ setting where synthetic data is used as input. When coarse quantization is applied without sufficiently granular scaling, distortion tends to accumulate in specific layers and ultimately degrades the output layer’s activation representation. This suggests that insufficient preservation of intermediate activations can negatively affect the quality of final inference. The problem becomes more severe in zero-shot scenarios due to the distributional mismatch introduced by synthetic data. In contrast, GranQ mitigates this issue by applying fine-grained channel-wise scaling, resulting in consistently stable activation distributions throughout the network. This phenomenon is clearly demonstrated in Figure 7, where GranQ maintains stable activations with noticeably reduced distortion compared to conventional quantization methods.
(a) Layer-wise Quantization
(b) GranQ (Ours)
Figure 7: Visualization of activation magnitudes across layers in ResNet-20. (a) illustrates the activation magnitudes with layer-wise quantization, whereas (b) presents the results after applying the proposed GranQ method. In each subplot, the left side depicts activation magnitudes before quantization (32-bit), whereas the right side shows them after quantization (3-bit). The proposed GranQ effectively preserves the activation distribution, mitigating distortions caused by quantization.