### v5

# GranQ: Granular Zero-Shot Quantization with Channel-Wise Activation Scaling in QAT

Abstract
Zero-shot quantization (ZSQ) enables neural network compression without original training data, making it a promising solution for restricted data access scenarios. To compensate for the lack of data, recent ZSQ methods typically rely on synthetic inputs generated from the full-precision model. However, these synthetic inputs often lead to activation distortion, especially under low-bit settings. To mitigate this, existing methods typically employ per-channel scaling, but they still struggle due to the severe computational overhead during the accumulation process. To overcome this critical bottleneck, we propose GranQ, a novel activation quantization framework that introduces an efficient pre-scaling strategy. Unlike conventional channel-wise methods that repeatedly perform scaling operations during accumulation, GranQ applies scaling factors in a pre-scaling step through fully vectorized computation, eliminating runtime scaling overhead. This design enables GranQ to maintain fine-grained quantization accuracy while significantly reducing computational burden, particularly in low-bit quantization settings. Extensive experiments under quantization-aware training (QAT) settings demonstrate that GranQ consistently outperforms state-of-the-art ZSQ methods across CIFAR and ImageNet. In particular, our method achieves up to 5.45% higher accuracy in the 3-bit setting on CIFAR-100 and even surpasses the full-precision baseline on CIFAR-10. Furthermore, GranQ achieves significant speedup in quantization latency over conventional per-channel methods, demonstrating improved efficiency. With these findings, we anticipate that GranQ will inspire future research beyond conventional ZSQ approaches centered on data generation and model fine-tuning. The official code is available at https://github.com/anonymus-orange/GranQ.
Introduction
Neural network compression has been extensively studied for the practical deployment of large-scale deep learning (DL) models. In particular, reducing model size while minimizing performance degradation is crucial for utilizing DL on edge devices (e.g., mobile phones, embedded systems, and drones). Major approaches to model compression include quantization (Hubara et al. 2018; Li et al. 2021; Gholami et al. 2022), pruning (LeCun, Denker, and Solla 1989; Frankle and Carbin 2018; Liu et al. 2018; Cheng, Zhang, and Shi 2024), knowledge distillation (Hinton 2015; Heo et al. 2019; Wang and Yoon 2021), and neural architecture search (Zoph 2016; Elsken, Metzen, and Hutter 2019), as surveyed in (Cheng et al. 2018; Deng et al. 2020). Among these, quantization has emerged as the most actively studied technique. It serves as an effective compression method by reducing unnecessary representational ranges in the model. However, it often requires fine-tuning or calibration to match full-precision (FP) model performance (Gholami et al. 2022). To address this, zero-shot quantization (ZSQ), also known as data-free quantization, has been proposed to compress models without the original training data.
Data Generation
(PTQ, QAT)
Calibration
Fine-tuning
ZeroQ (CVPR 20)
GDFQ (ECCV 20)
DSG (CVPR 21)
MixMix (CVPR 21)
Qimera (NeurIPS 21)
IntraQ (CVPR 22)
GENIE (CVPR 23)
AdaDFQ (CVPR 23)
Casual-DFQ (ICCV 23)
TexQ (NeurIPS 23)
RIS (AAAI 24)
GenQ (ECCV 24)
SQuant (ICLR 22)
UDFC (ICCV 23)
AIT (CVPR 22)
PLF (CVPR 24)
SynQ (ICLR 25)
GranQ (Ours)
Table 1: Categorizing ZSQ algorithms. PTQ denotes post-training quantization, and QAT denotes quantization-aware training. Data generation methods typically support both PTQ and QAT, whereas violet-marked methods utilize fine-tuning for data generation and support only QAT.
Since the introduction of ZeroQ (Cai et al. 2020), studies on ZSQ have advanced in two main directions. The first direction focuses on data generation, where synthetic data are created from the FP model. The second direction focuses on effectively applying the activation distributions of the synthetic data to the quantized (Q) model. This second direction is further divided into post-training quantization (PTQ), which calibrates the activation distributions, and quantization-aware training (QAT), which fine-tunes the Q model directly. We categorize existing ZSQ methods based on these directions, as summarized in Table 1.
(a) Layer-wise Quantization
(b) GranQ (Channel-wise, Ours)
Figure 1: Comparison between (a) layer-wise quantization and (b) GranQ on the CIFAR-10. Each subfigure visualizes the 32-bit FP (left) and 3-bit quantized (right) activations of the first ResNet-20 layer. GranQ better preserves the original activation with minimal distortion.
First, data generation studies focus on generating high-quality data to effectively train the Q model (Xu et al. 2020; Qian et al. 2023; Li et al. 2024). Meanwhile, calibration (PTQ) studies aim to minimize the quantization error by calibrating the Q model without additional training (Guo et al. 2022; Bai et al. 2023). Finally, fine-tuning (QAT) studies focus on transferring key information from the FP model to the Q model through knowledge distillation (Choi et al. 2022; Fan et al. 2024; Hong et al. 2025; Kim, Kim, and Kang 2025).
However, despite extensive studies, severe performance degradation in low-bit quantization remains unresolved. To address this issue, we performed an in-depth analysis of the ZSQ process, focusing on why low-bitwidth settings still suffer from performance loss even after QAT fine-tuning. Our findings reveal that quantization errors mainly stem from the loss of activation values instead of data quality or training methods. Notably, we found that layer-wise (per-tensor) quantization is no longer suitable for preserving activations in ZSQ, as it leads to coarse and inaccurate representations.
Based on this analysis, we introduce GranQ, a novel ZSQ method that achieves efficient per-channel quantization via vectorized pre-scaling of input-dependent activations. This dynamic adjustment minimizes activation loss and preserves the original activation values by reducing quantization errors, as shown in Figure 1. The proposed method effectively handles activation loss in low-bit quantization and achieves state-of-the-art (SOTA) performance in QAT settings on the CIFAR and ImageNet datasets. Furthermore, we apply vectorization to the pre-scaling step, which is typically omitted in conventional channel-wise quantization, thereby reducing latency and enabling fine-grained activation quantization. Our contributions can be summarized as follows:
  • We identify critical limitations of layer-wise activation quantization in low-bit ZSQ. Our findings reveal that conventional activation quantization methods relying on layer-wise granularity suffer from significant activation loss. This limitation becomes more severe in ZSQ settings with synthetic data.
  • We propose GranQ, a novel method that supports granular quantization and maintains computational efficiency. Although per-channel activation quantization is known to improve precision, it has not been widely adopted due to its high computational cost. We address this by introducing vectorized pre-scaling, which integrates per-channel scaling into the quantization step, allowing accumulation to proceed efficiently without runtime scaling overhead. To the best of our knowledge, our approach is the first to address the ZSQ problem.
  • We achieve SOTA performance over existing ZSQ methods through extensive evaluation. Specifically, on the CIFAR-100 dataset, in the 3-bit quantization setting, GranQ achieves an accuracy of 62.73%, improving by 5.45% over the latest method on ResNet-20. Furthermore, on the CIFAR-10 dataset in the 5-bit quantization setting, GranQ achieves an accuracy of 94.06%, slightly exceeding the FP model performance by 0.17%.
Related Work
Quantization
Quantization reduces the representational range of deep neural networks (DNNs) and minimizes memory usage. It is typically categorized into post-training quantization (PTQ) and quantization-aware training (QAT) (Gholami et al. 2022; Banner, Nahshan, and Soudry 2019). PTQ applies quantization after training without further updates, which makes it efficient and easy to implement. However, it is sensitive to scaling errors and often relies on calibration to improve accuracy. In contrast, QAT incorporates quantization during training and optimizes quantized activations using the straight-through estimator (Gholami et al. 2022; Yin et al. 2019). Both methods require data. PTQ uses it for calibration, while QAT uses it for fine-tuning.. In practice, access to training data is often limited. To address this, zero-shot quantization (ZSQ) has been proposed to perform quantization without using original data (Xu et al. 2020).
Zero-shot Quantization
As summarized in Table 1, ZSQ was initially introduced by ZeroQ (Cai et al. 2020), leading to the development of various ZSQ algorithms. These studies have explored diverse approaches to improve quantization performance under data-free settings. Recent studies have focused on generating high-quality synthetic data (Jeon, Lee, and Kim 2023; Li et al. 2024) and developing more effective methods to utilize them (Choi et al. 2022; Hong et al. 2025; Kim, Kim, and Kang 2025). However, current ZSQ methods heavily rely on augmented synthetic inputs. This causes large variation in activation scales across channels. In practice, such input-dependent channel-wise quantization is rarely adopted due to its high computational cost. Therefore, recent studies have applied channel-wise activation quantization either in a limited conditions within QAT settings (Cho and Yoo 2020), or exclusively in PTQ scenarios with access to a small amount of data (Yvinec et al. 2023; Xu et al. 2024). This highlights the importance of efficient channel-wise activation quantization in QAT.
Preliminaries and Problem Definition
In this section, we provide a new definition of the existing ZSQ problem and introduce the preliminaries. The detailed mathematical notations are listed in Table 6 in the appendix.
Activation Quantization
Activation quantization reduces the precision of intermediate activation values by converting them into low-bitwidth integers (Choi et al. 2018). It is commonly used with weight quantization to compress models, and the linear quantization scheme is most widely adopted. This process typically consists of quantization, scaling parameter computation, and dequantization (Hubara et al. 2018; Gholami et al. 2022; Fang et al. 2020).
The quantization operator $\mathcal{Q}$ maps a floating-point value $x$ into an integer $x_{q}$ using a scaling factor $s$ and zero-point $z$:
$x_{q}=\mathcal{Q}(x,s,z)=\left\lfloor\frac{x}{s}+z\right\rceil$ (1)
Here, $s$ normalizes the activation range, and $z$ shifts the scaled value to account for asymmetric quantization. These parameters are computed as:
$s=\frac{x_{\max}-x_{\min}}{2^{b}-1},\quad z=\left\lfloor\frac{-x_{\min}}{s}\right\rceil$ (2)
where $x_{\min}$ and $x_{\max}$ denote the range of activation values, and $b$ is the bit-width. A large range yields a large $s$, increasing quantization error, a smaller range allows finer resolution. As shown below, in quantized models, per-channel scaling is embedded directly into the accumulation path, requiring each channel to be individually scaled during computation:
$y_{l}=\sum_{c=1}^{C}w_{c}\cdot s_{c}\cdot(x_{q,c}-z_{c})$ (3)
However, applying the per-channel scaling factor $s_{c}$ inside the accumulation loop introduces computational overhead, making efficient integer-domain execution difficult. This scaling bottleneck hinders parallelism in core operations such as convolution and matrix multiplication, reducing the efficiency of quantized models (Nagel et al. 2021).
Problem Definition
Existing activation quantization methods are designed based on layer-wise quantization to minimize computational cost. However, a fixed scaling factor struggles to handle varying activations. This issue becomes more severe under low-bitwidth settings in ZSQ environments. The following outlines the primary challenges faced in ZSQ under low-bitwidth settings.
(Problem 1) Coarse Activation Quantization by Single-Range Scaling
Conventional activation quantization uses a single scaling range per layer for efficiency, which was acceptable at higher bitwidths or when real data is available. However, this design becomes problematic in ZSQ, where range adjustment relies on synthetic data. These synthetic distributions are often biased or irregular, making precise quantization more difficult.
(Problem 2) Accumulation Bottleneck in Channel-wise Quantization The main challenge in channel-wise quantization lies not in quantizing activations but in the repeated per-channel scaling during accumulation. This disrupts vectorized execution and significantly increases computational overhead.
Observation and Methodology
We observe the causes of activation distortion during quantization and analyze how to mitigate them. Based on this, we introduce our method, GranQ.
Observation
Activaiton Quantization Layer-wise Channel-wise (GranQ)
$Avg.\frac{X\cdot Q}{\|X\|_{2}\|Q\|_{2}}(\uparrow)$
$Avg.\frac{\|X-Q\|_{2}}{\|X\|_{2}}(\downarrow)$
0.5111
0.3129
0.6835
0.1063
Table 2: Average cosine similarity and relative error for 3-bit activation quantization on ResNet-20 (CIFAR-100). GranQ shows lower activation distortion than layer-wise, with higher similarity and lower error.
[…]
Furthermore, Figure 1(a) illustrates how layer-wise quantization leads to coarse activation representations across layers. Fine-grained activation values are often mapped to limited discrete levels, resulting in a loss of detail and reduced representational capacity. This degradation contributes to decreased inference performance.
Methodology
Based on our analysis, we propose GranQ, a fine-grained ZSQ method designed to reduce computational overhead, as illustrated in Figure 2. GranQ applies channel-wise scaling through activation decomposition, and achieves fast parallel processing via pre-scaling for parallel accumulation, which enables efficient computation of both scaling and quantization.
Algorithm 1 GranQ: Pre-scaling for Parallel Accumulation
1:Activation tensors $\{A_{l}\}_{l=1}^{L}$, where $A_{l}\in\mathbb{R}^{C\times H\times W}$, quantization bit $b$, weight tensors $\{w_{l,c}\}_{l=1,c=1}^{L,C}$
2:Output vectors $\{y_{l}\}_{l=1}^{L}$
3:for each layer $l$ in $L$ do
4:  $A_{l}\leftarrow\text{reshape}(A_{l})\in\mathbb{R}^{C\times(HW)}$
5:  $\vec{A}_{\min},\vec{A}_{\max}\leftarrow\min_{h,w}(A_{l}),\max_{h,w}(A_{l})$
6:  $\vec{s}_{l}\leftarrow(\vec{A}_{\max}-\vec{A}_{\min})/(2^{b}-1)$
7:  $\vec{z}_{l}\leftarrow\left\lfloor-\vec{A}_{\min}/\vec{s}_{l}\right\rceil$
8:  $\vec{A}_{q,l}\leftarrow\left\lfloor A_{l}/\vec{s}_{l}+\vec{z}_{l}\right\rceil$ $\triangleright$ quantize (int mapping)
9:  $\vec{\tilde{A}}_{l,c}\leftarrow\vec{s}_{l}\cdot(\vec{A}_{q,l}-\vec{z}_{l})$ $\triangleright$ dequantize & pre-scale
10:  Compute output:
$y_{l}=\sum_{c=1}^{C}w_{l,c}\cdot\vec{\tilde{A}}_{l,c}$
11:end for
12:return $\{y_{l}\}_{l=1}^{L}$
Activation Decomposition
We propose activation decomposition, which reshapes each activation map from a three-dimensional tensor of shape $(C\times H\times W)$ into a two-dimensional matrix of shape $(C\times HW)$. This transformation enables decomposition of the activation map along the channel axis, allowing each channel to be processed independently. It also facilitates vectorized computation of per-channel statistics (e.g., min and max) required for scaling factor calculation during quantization. We define the activation decomposition as shown in Equation 4.
$\begin{gathered}A_{l}=\left[A_{l}(1,:,:),\,A_{l}(2,:,:),\,\dots,\,A_{l}(C,:,:)\right]\in\mathbb{R}^{C\times H\times W}\\ \vec{A}_{\min}=\left[\min_{(h,w)}A_{l}(c,h,w)\right]_{c=1}^{C}\in\mathbb{R}^{C}\\ \vec{A}_{\max}=\left[\max_{(h,w)}A_{l}(c,h,w)\right]_{c=1}^{C}\in\mathbb{R}^{C}\end{gathered}$ (4)
[…]
Pre-scaling for Parallel Accumulation
$\vec{s}=\frac{\vec{A}_{\max}-\vec{A}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{A}_{\min}}{\vec{s}}\right\rceil$ (5)
$y_{l}=\sum_{c=1}^{C}w_{l,c}\cdot\vec{\tilde{A_{l,c}}},\quad\text{where }\vec{\tilde{A_{l,c}}}=\vec{s}_{c}\cdot\left(\left\lfloor\frac{\vec{A}_{l,c}}{\vec{s}_{c}}+\vec{z}_{c}\right\rceil-\vec{z}_{c}\right)$ (6)
In the pre-scaling for parallel accumulation stage, the scaling factors $\vec{s}_{c}$ and zero-points $\vec{z}_{c}$ are computed in a vectorized form across channels to dynamically adapt to the diverse activation distributions induced by synthetic inputs. This contrasts with conventional per-channel quantization, which applies fixed scaling parameters per channel regardless of input variation.
As shown in Equation 6, each activation $\vec{A}_{l,c}$ is first quantized using its corresponding vectorized scaling and zero-point, and then dequantized to reconstruct the floating-point value. The resulting $\vec{\tilde{A_{l,c}}}$ represents a pre-scaled activation that is already adjusted for accumulation. By applying vectorized scaling, quantization, and dequantization uniformly across channels, this process enables efficient integer-domain computation and parallel accumulation with weights $w_{l,c}$.
Experiments
In this section, we thoroughly evaluate the effectiveness of GranQ. Experiments are conducted on diverse benchmark datasets, with the performance compared with those of existing ZSQ methods.
Experimental Setup and Details
The experiments were conducted using widely adopted ZSQ evaluation datasets, including CIFAR-10, CIFAR-100 (Krizhevsky, Hinton et al. 2009), and ImageNet (ILSVRC 2012) (Deng et al. 2009) validation datasets. For the CIFAR datasets, ResNet-20 (He et al. 2016) was used as the quantization model, whereas ResNet-18 (He et al. 2016), ResNet-50 (He et al. 2016), and MobileNetV2 (Sandler et al. 2018) were employed for ImageNet. All experiments were conducted using the SGD optimizer (Ruder 2016) with a momentum of 0.9 and weight decay of 1e-4. The CIFAR-10 and CIFAR-100 experiments were each conducted for 200 epochs, with batch sizes of 16 and 200, respectively. For ImageNet, we trained for 400 epochs with a batch size of 16. The initial learning rate was set to 1e-4 for CIFAR-10 and CIFAR-100, and 1e-5 for ImageNet, with multi-step learning rate decay applied. The decay steps were set to 100, 200, and 300 epochs for CIFAR, and at 350 and 400 epochs for ImageNet, with a decay rate of 0.1. We compared our method with existing ZSQ methods (Xu et al. 2020; Choi et al. 2022; Qian et al. 2023; Chen et al. 2024; Bai et al. 2024; Fan et al. 2024; Li et al. 2024; Hong et al. 2025; Kim, Kim, and Kang 2025). For data generation, we followed the AdaDFQ (Qian et al. 2023) approach based on ACGAN (Odena, Olah, and Shlens 2017). Further implementation details are provided in the appendix, Section A and A. In our ablation study, layer-wise quantization was applied to all layers containing activation functions, while channel-wise quantization was performed per channel at the batch level.
Dataset
Model
(FP 32)
Bits
GDFQ
(ECCV 20)
ARC+AIT
(CVPR 22)
AdaDFQ
(CVPR 23)
TexQ
(NeurIPS 24)
AIT+RIS
(AAAI 24)
GenQ
(ECCV 24)
AKT
(SAC 25)
SynQ
(ICLR 25)
GranQ
Cifar-10
ResNet-20
3ww3aa 75.11 - 84.89 86.47 - - 86.76 88.11 91.37
4ww4aa 90.25 90.49 92.31 92.68 92.59 - 92.64 92.76 93.52
5ww5aa 93.38 92.98 93.81 - 93.59 - 93.83 - 94.06
Cifar-100
ResNet-20
3ww3aa 47.61 - 52.74 55.87 - - 54.68 57.28 62.73
4ww4aa 63.39 61.05 66.81 67.18 65.99 - 66.94 67.34 68.79
5ww5aa 66.12 68.40 69.93 - 69.55 - 69.75 - 70.05
ImageNet
ResNet-18
3ww3aa 20.23 - 38.10 50.28 - 68.18 49.88 52.02 64.41
4ww4aa 60.60 65.73 66.53 67.73 67.55 70.03 65.89 67.90 70.39
5ww5aa 68.49 70.28 70.29 - 70.59 - 69.40 - 71.31
MobileNetV2
3ww3aa 1.46 - 28.99 32.80 - 59.15 30.56 34.21 62.42
4ww4aa 59.43 66.47 65.41 67.07 - 69.65 64.85 67.27 70.62
5ww5aa 68.11 71.96 71.61 - - - 71.71 - 72.49
ResNet-50
3ww3aa 0.31 - 17.63 25.27 - 73.99 24.50 26.89 70.76
4ww4aa 54.16 68.27 68.38 70.72 71.54 76.10 68.75 71.05 76.63
5ww5aa 71.63 76.00 76.03 - 76.36 - 75.90 - 77.58
Table 3: Accuracy Evaluation of QAT Methods for ZSQ. $w$ and $a$ represent weight and activation, respectively. Bold values indicate the best accuracy, and underlined values denote the second-best accuracy. $\dagger$ indicates our re-implementation with the official code.
CIFAR-100 ResNet-20 (70.33%)
Method GDFQ Qimera+AIT AdaDFQ AdaDFQ+AKT
Baseline +GranQ Baseline +GranQ Baseline +GranQ Baseline +GranQ
3ww3aa 47.61 59.04+11.43 45.70 60.42+14.72 52.74 62.73+9.99 54.68 62.01+7.33
4ww4aa 63.39 66.97+3.58 65.80 68.08+2.28 66.81 68.79+1.98 66.94 68.77+1.83
5ww5aa 66.12 68.96+2.84 69.26 70.14+0.88 69.93 70.05+0.12 69.75 70.21+0.46
Table 4: Accuracy of existing SOTA methods with the integration of GranQ. GDFQ and AdaDFQ focus on data generation, whereas Qimera+AIT and AdaDFQ+AKT primarily enhance quantized model training.
Figure 3: Latency of ResNet-20 quantization across batch sizes on CIFAR-100 with 3-bit setting.
Figure 4: Relative quantization error across layers in ResNet-20 with 3-bit quantization on CIFAR-100.
Performance Evaluation
We evaluated the performance of GranQ against SOTA ZSQ methods, with results summarized in Table 3. All comparison experiments were conducted under 3, 4, and 5-bit quantization settings.
CIFAR-10/100. As indicated in Table 3, GranQ consistently achieved the highest accuracy across all bitwidths. For CIFAR-10, it attained accuracies of 94.06% (5-bit), 93.52% (4-bit), and 91.37% (3-bit). For CIFAR-100, the results were 70.05% (5-bit), 68.79% (4-bit), and 62.73% (3-bit). Notably, GranQ outperformed SynQ (Kim, Kim, and Kang 2025), the previous SOTA, by +5.45% in the CIFAR-100 3-bit setting. This result demonstrates its ability to effectively overcome the limitations of conventional low-bitwidth quantization techniques. Overall, GranQ consistently outperformed existing methods across all bitwidths in both CIFAR-10 and CIFAR-100, with particularly strong improvements in the 3-bit quantization setting. Remarkably, in the CIFAR-10 5-bit setting, GranQ even surpassed the performance of the FP model. These results suggest that we can effectively apply GranQ to small- and medium-scale datasets.
ImageNet. In the ImageNet experiments, GranQ achieved competitive performance across various bitwidths. Specifically, for ResNet-18, it attained top accuracies of 70.39% (4-bit) and 71.31% (5-bit). For ResNet-50, it achieved the highest accuracies of 76.63% (4-bit) and 71.31% (5-bit). Additionally, in the MobileNetV2 setting, GranQ achieved SOTA performance across all bitwidths (3, 4, and 5-bit). In the 3-bit setting, GranQ achieved the second-best accuracies, with 64.41% on ResNet-18 and 70.76% on ResNet-50. These results are -3.77% and -3.23% lower than those of GenQ, respectively. However, GenQ uses a pre-trained diffusion model (Ho, Jain, and Abbeel 2020) and prompts during data generation, leading to slower speeds compared with those of typical GAN-based methods (Xu et al. 2020). In contrast, GranQ enhances the quantization mechanism without heavily relying on diffusion-based data generation, as GenQ does. Although GenQ demonstrates strong performance, its code is not publicly available, and thus we could not include it in our experiments. Nevertheless, since our method is compatible with existing QAT frameworks, it could potentially benefit from GenQ’s data generation if integration becomes possible.
Ablation Study
Effectiveness Evaluation
GranQ consistently demonstrates performance improvements when applied to various ZSQ methods. As summarized in Table 4, GranQ achieves steady performance improvements when integrated with existing SOTA ZSQ techniques.
First, we analyzed the impact of GranQ on data synthesis-based quantization methods, specifically GDFQ (Xu et al. 2020) and AdaDFQ (Qian et al. 2023). The integration of GranQ into GDFQ (Xu et al. 2020) led to a notable improvement, with accuracy rising from 47.61% to 59.04% (+11.43%) in the 3-bit setting and from 63.39% to 66.97% (+3.58%) in the 4-bit setting. Similarly, AdaDFQ (Qian et al. 2023) exhibited improvements of +9.99% in the 3-bit setting (from 52.74% to 62.73%) and +1.98% in the 4-bit setting (from 66.81% to 68.79%). These results suggest that GranQ effectively reduces quantization errors when combined with data synthesis-based quantization methods, leading to enhanced model performance.
Furthermore, GranQ also exhibits significant improvements in methods focused on training quantized models, such as Qimera+AIT (Choi et al. 2022) and AdaDFQ+AKT (Hong et al. 2025). For Qimera+AIT (Choi et al. 2022), the 3-bit accuracy increased from 45.70% to 60.42% (+14.72%). Similarly, for AdaDFQ+AKT (Hong et al. 2025), the accuracy improved from 54.68% to 62.01% (+7.33%). These findings demonstrate that GranQ is not only effective in data synthesis-based methods but also enhances performance during the model training process.
Efficiency Evaluation
Method
Quantization
Pre-Scaling
Accuracy (%)
Quantization
Latency (ms)
Training
Time (sec/epoch)
Layer-wise 49.98 0.227 10.92
Channel-wise (Baseline) 62.68 103.671 17370.54
Channel-wise (GranQ) 62.73 0.236 12.43
Table 5: Ablation study on the impact of pre-scaling and vectorization in quantization. ✓indicates whether vectorization is applied. Latency is measured with a batch size of 16 using AdaDFQ on ResNet-20 (CIFAR-100, 3-bit).
While fine-grained activation quantization often incurs high computational cost due to per-channel scaling, GranQ maintains efficiency by vectorizing the scaling computation. By reshaping activation tensors, our method eliminates the overhead of iterative channel-wise operations. To assess this in practice, we measured quantization latency across various batch sizes (16, 32, 64, 128, and 200). Figure 4 compares three methods: layer-wise quantization, channel-wise quantization with scalar-loop scaling (Algorithm 2 in appendix), and our method (GranQ), which efficiently parallelizes scaling across channels.
The two channel-wise methods perform per-channel quantization. The conventional method computes scaling factors sequentially, whereas GranQ pre-applies them in a vectorized manner across channels, enabling pre-scaling for efficient parallel accumulation. A key observation from Figure 4 is that GranQ achieves substantial accuracy improvement with only a minimal latency overhead compared to conventional layer-wise quantization. Furthermore, Table 2 and Figure 4 shows that GranQ substantially reduces quantization error and more effectively preserves activation information. These results demonstrate that GranQ enables fine-grained quantization with significantly improved efficiency.
Additionally, GranQ demonstrates that per-channel quantization can be efficiently executed by vectorizing scaling factor computation. As shown in Table 5, it achieves a quantization latency of 0.236 ms, comparable to the layer-wise baseline (0.227 ms), and significantly outperforms the conventional channel-wise method (103.671 ms) suffering from scalar-loop overhead.
Discussion
Why GranQ is Effective?
In layer-wise quantization, representing the entire activation range with a single scaling factor makes it difficult to reflect fine-grained distribution changes. These limitations become more severe in low-bit settings, especially when synthetic data exhibit distributional shifts. To overcome these issues, GranQ pre-computes per-channel scaling in a vectorized form, enabling efficient parallel accumulation. By integrating scaling into the quantization phase, we removed runtime scaling during accumulation, allowing precise and efficient parallel execution.
Limitation and Future-work
In ZSQ, synthetic data often exhibit broader and noisier activation distributions (see Figure 6 in appendix). While GranQ effectively handles this in QAT through dynamic per-channel scaling, its static form is less suited for PTQ. Future work will explore adapting GranQ to better support PTQ.
Conclusion
Layer-wise and channel-wise quantization are both commonly used in weight quantization, as weights remain fixed after training. In contrast, it is harder to apply channel-wise activation quantization because input-dependent activations make fine-grained scaling difficult. These issues are especially critical in ZSQ, with data synthesized from FP statistics. In such scenarios, synthetic activations often deviate from real data, amplifying quantization error and complicating accurate scaling. In addition, channel-wise quantization poses a computational burden due to the need for per-channel scaling during accumulation. Since scaling factors differ across channels, accumulation requires separate multiplications, limiting parallelism and increasing latency. This overhead discourages the use of channel-wise methods in practice, despite their higher representational fidelity. To address these issues, we propose GranQ, which vectorizes scaling to enable fine-grained quantization and efficient accumulation across channels. Through extensive experiments, we demonstrate that GranQ achieves notable accuracy and efficiency. Our study highlights the importance of activation preservation in ZSQ and proposes GranQ as an insightful direction to address this challenge.
Appendix A Appendix
Notation
We provide an overview of the main symbols and definitions used throughout the paper, along with their precise mathematical formulations.
Symbol Definition
AA Original activation tensor
$\vec{A}_{q}$ Quantized activation vector (per channel)
$\vec{\tilde{A}}$ Pre-scaled activation
$\vec{A}_{\min},\vec{A}_{\max}$ Per-channel min and max activation values
$\vec{s}$ Per-channel scaling factor vector
$\vec{z}$ Per-channel zero-point vector
$b$ Number of quantization bits
$C,H,W$ Channel, Height, Width dimensions of activation
$\mathbb{R}^{C\times H\times W}$ Activation tensor space
$\lfloor\cdot\rceil$ Rounding function
$\cdot$ Element-wise multiplication
$\sum_{c=1}^{C}$ Channel-wise summation
$w_{c}$ Weight for channel $c$
$y_{l}$ Output after accumulation at layer $l$
Mathematical Expressions
Scaling Factor $\vec{s}=\frac{\vec{A}_{\max}-\vec{A}_{\min}}{2^{b}-1}$
Zero-point $\vec{z}=\left\lfloor\frac{-\vec{A}_{\min}}{\vec{s}}\right\rceil$
Forward Quantization $\vec{A}_{q}=\left\lfloor\frac{A}{\vec{s}}+\vec{z}\right\rceil$
Pre-scaled Dequantization $\vec{\tilde{A}}=\vec{s}\cdot(\vec{A}_{q}-\vec{z})$
Output Accumulation $y=\sum_{c=1}^{C}w_{c}\cdot\vec{\tilde{A}}_{c}$
Cosine Similarity $\frac{X\cdot Q}{\|X\|_{2}\|Q\|_{2}}$
Relative Error $\frac{\|X-Q\|_{2}}{\|X\|_{2}}$
Table 6: Notation and expressions for vectorized channel-wise activation quantization with pre-scaling.
Details of Experimental Setup
We describe the datasets, experimental setup, and baselines to ensure reproducibility and highlight both the effectiveness and implementation accessibility of our approach.
Datasets
This paper outlines the experimental settings for CIFAR-10 (Krizhevsky, Hinton et al. 2009), CIFAR-100 (Krizhevsky, Hinton et al. 2009), and ImageNet (ILSVRC-2012) (Deng et al. 2009) datasets. CIFAR-10 and CIFAR-100 contain 10 and 100 classes, respectively, with 32 × 32 resolution color images, and both have 50 K training and 10 K test images. ImageNet comprises 1,000 classes, with 1.2 M training and 50 K validation images, all resized to 224 × 224 for our experiments. We employ these widely used benchmarks for zero-shot quantization (ZSQ) to assess and compare the performance of GranQ against those of existing approaches.
Experimental Setup
The data generation process followed the settings of AdaDFQ (Qian et al. 2023). Specifically, ACGAN (Odena, Olah, and Shlens 2017) was used for data generation with the Adam optimizer, set with a momentum of 0.9 and a learning rate of 1e-3. All hyper-parameters such as $\alpha_{ds}$, $\alpha_{as}$, $\lambda_{l}$, and $\lambda_{u}$ followed the default settings from AdaDFQ (Qian et al. 2023). For the fine-tuning process, a warmup phase was applied to ensure training stability, with 4 epochs out of 200 for CIFAR and 50 epochs out of 400 for ImageNet. The batch size was set to 16 for CIFAR-10 and ImageNet, and 200 for CIFAR-100. All experiments used Nesterov-based SGD (Nesterov 1983) as the optimizer, with momentum 0.9 and weight decay 1e-4. In CIFAR experiments, the weight decay of 1e-4 was reduced by a factor of 0.1 every 100 epochs, whereas for ImageNet, decay was applied at 350 and 400 epochs (for ResNet-50 at 150, 300, and 400 epochs). GranQ was implemented using PyTorch (Paszke et al. 2019), and visualizations were generated with matplotlib (Hunter 2007). All experiments were conducted on an NVIDIA RTX 3090 GPU.
Comparison Methods
GranQ is a quantization-aware training (QAT) method for zero-shot quantization (ZSQ). Therefore, to evaluate the proposed approach, we conducted comparative experiments using the following state-of-the-art methods:
  • GDFQ (Xu et al. 2020): GDFQ (Generative low-bit data-free quantization) is the first approach to train quantized models using synthetic data through knowledge distillation. It demonstrated that effective quantization is possible even without access to original data.
  • AIT (Choi et al. 2022): AIT was the first to show the negative impact of cross-entropy loss when training quantized models with distillation. It also provided a systematic analysis of training strategies in the ZSQ domain.
  • AdaDFQ (Qian et al. 2023): AdaDFQ (Adaptive data-free quantization) highlighted the generalization issues of generated data under various quantization settings and proposed a more adaptable data generation method.
  • TexQ (Chen et al. 2023): TexQ introduced a data generation method that preserves texture distributions, thereby enhancing the quality of generated data in ZSQ environments.
  • RIS (Bai et al. 2024): RIS (Robustness-guided image synthesis) addressed the low semantic quality of generated data and proposed using perturbations to improve data diversity.
  • GenQ (Li et al. 2024): GenQ introduced the use of stable diffusion in ZSQ, enabling the generation of high-resolution, realistic data.
  • AKT (Hong et al. 2024): AKT (Advanced knowledge transfer) proposed refining feature information during distillation to improve the training of quantized models.
  • SynQ (Kim, Kim, and Kang 2025): SynQ enhanced the training of quantized models by reducing noise in generated data using low-pass filtering and aligning activation maps.
Algorithm 2 Conventional Channel-Wise Quantization (Scalar-loop for scaling)
1:Activation tensors $\{A_{l}\}_{l=1}^{L}$, where $A_{l}\in\mathbb{R}^{C\times H\times W}$, quantization bit $b$, weights $\{W_{l}\}_{l=1}^{L}$
2:Output vectors $\{y_{l}\}_{l=1}^{L}$
3:for each layer $l$ in $L$ do
4:  $A_{l}\leftarrow\text{reshape}(A_{l})\in\mathbb{R}^{C\times(HW)}$
5:  for each channel $c$ in $C$ do
6:   $s_{c}=\frac{\max(A_{l,c})-\min(A_{l,c})}{2^{b}-1}$
7:   $z_{c}=\left\lfloor-\frac{\min(A_{l,c})}{s_{c}}\right\rceil$
8:   $A_{q,l}[c,:]=\left\lfloor\frac{A_{l}[c,:]}{s_{c}}+z_{c}\right\rceil$
9:  end for
10:  Compute output:
$y_{l}=\sum_{c=1}^{C}w_{l,c}\cdot\left(s_{c}\cdot(A_{q,l}[c,:]-z_{c})\right)$
11:end for
12:return $\{y_{l}\}_{l=1}^{L}$
Algorithm
While prior works apply vectorized computation only to the quantization step, they compute per-channel scaling parameters through iterative operations. GranQ also vectorizes this scaling step, enabling fully parallel quantization. Algorithm 1 summarizes the process with pre-scaling.
In Algorithm 1, the entire quantization procedure is implemented using vectorized tensor operations across all channels, enabling efficient accumulation without runtime scaling. Specifically, the per-channel minimum and maximum values are computed in parallel, and the corresponding scaling factors and zero-points are applied via broadcasting. This removes the need for channel-wise loops and enables efficient accumulation without runtime scaling, making the procedure highly suitable for modern hardware accelerators.
In contrast, Algorithm 2 adopts a conventional channel-wise quantization scheme, where scaling parameters are computed iteratively for each channel using a for-loop. Although the quantization and dequantization steps are applied in a vectorized manner per channel, the overall process remains sequential due to scalar-based scaling. This incurs significant computational overhead, particularly in networks with a large number of channels.
By removing this bottleneck, GranQ achieves both higher parallelism and lower latency, which are particularly beneficial in low-bitwidth quantization scenarios.
Extra Analysis
Data Distribution Analysis
(a) Original inputs (Skewness: 0.2599)
(b) Synthetic inputs (Skewness: 0.4877)
Figure 5: Pixel distribution comparison between (a) original and (b) synthetic inputs on ImageNet. Synthetic inputs generated by AdaDFQ (Qian et al. 2023) exhibit a rightward shift with higher skewness, indicating more extreme activations and highlighting the need for granular scaling.
Figure 5 compares the pixel distributions of original and synthetic inputs derived from ImageNet. The skewness of the original input is 0.2599, whereas that of the synthetic input increases to 0.4877, with the distribution shifted toward higher pixel values. This indicates that synthetic data, often generated without real image priors, tends to contain a greater number of extreme activation values. As a result, the likelihood of outliers increases, which fixed-range quantization cannot effectively handle. In particular, when a single global range is applied across all channels, the presence of outliers can compress the majority of values into a narrow dynamic range, leading to severe quantization distortion. These observations underscore the limitations of conventional scaling strategies.
Analysis of Activation Preservation
We analyzed how activation distortion propagates across layers in a ZSQ setting where synthetic data is used as input. When coarse quantization is applied without sufficiently granular scaling, distortion tends to accumulate in specific layers and ultimately degrades the output layer’s activation representation. This suggests that insufficient preservation of intermediate activations can negatively affect the quality of final inference. The problem becomes more severe in zero-shot scenarios due to the distributional mismatch introduced by synthetic data. In contrast, GranQ mitigates this issue by applying fine-grained channel-wise scaling, resulting in consistently stable activation distributions throughout the network. This phenomenon is clearly demonstrated in Figure 6, where GranQ maintains stable activations with noticeably reduced distortion compared to conventional quantization methods.
(a) Layer-wise Quantization
(b) GranQ (Ours)
Figure 6: Visualization of activation magnitudes across layers in ResNet-20. (a) illustrates the activation magnitudes with layer-wise quantization, whereas (b) presents the results after applying the proposed GranQ method. In each subplot, the left side depicts activation magnitudes before quantization (32-bit), whereas the right side shows them after quantization (3-bit). The proposed GranQ effectively preserves the activation distribution, mitigating distortions caused by quantization.